### v1

# The Complexity of Maximizing the MST-ratio

Abstract
Given a finite set of red and blue points in ${{\mathbb{R}}}^{d}$, the MST-ratio is the combined length of the Euclidean minimum spanning trees of red points and of blue points divided by the length of the Euclidean minimum spanning tree of the union of them. The maximum MST-ratio of a point set is the maximum MST-ratio over all non-trivial colorings of its points by red and blue. We prove that the problem of finding the maximum MST-ratio of a given point set is NP-hard when the dimension is a part of the input. Moreover, we present a ${O}(n^{2})$ running time $3$-approximation algorithm for it. As a part of the proof, we show that in any metric space, the maximum MST-ratio is smaller than $3$. Additionally, we study the average MST-ratio over all colorings of a set of $n$ points. We show that this average is always at least $\frac{n-2}{n-1}$, and for $n$ random points uniformly distributed in a $d$-dimensional unit cube, the average tends to $d]{2}$ in expectation as $n$ goes to infinity.
1 Introduction
Recently, motivated by applications in spatial biology, Cultrera et al. [6] studied the interactions between color classes in a colored point set from a topological point of view. They introduced the concept of MST-ratio as one of the measures for mingling points with different colors in a colored point set. Investigating this measure opens the door to interesting discrete geometry questions related to Euclidean minimum spanning trees, or EMST for short, which is the main focus of this paper. For a set $P$ of $n$ points in the Euclidean space ${{\mathbb{R}}}^{d}$, the Euclidean minimum spanning tree of $P$, denoted by $EMST(P)$, is the minimum spanning tree of the complete geometric graph on $P$, where the edges have weights equal to the Euclidean distance between the two endpoints. For a partition $P=R\cup B$ of $P$ into red and blue points, the EMST-ratio (sometimes referred to as the MST-ratio) of $P$ and $R$ is the combined length of the Euclidean minimum spanning trees of $R$ and $B$, divided by the length of the Euclidean minimum spanning trees of $P$:
$\mu(P,R)=\frac{|EMST(R)|+|EMST(B)|}{|EMST(P)|}.$
How much longer can the EMSTs of two finite sets be compared to the EMST of their union? Given this union, $P$, we are interested in the maximum ratio, $\gamma(P)$, over all non-trivial partitions of $P$ into two sets. Upper and lower bounds for $\gamma$ of certain classes of point sets can be found in [7, 10], see Section 1.1 for further details. However, not a great deal of research has been conducted in the area of maximum EMST-ratio for higher dimensional Euclidean space.
Moreover, the question of whether there is an efficient algorithm to compute $\gamma$ for a given point set has yet been unanswered. The question can also be raised in the abstract setting. Namely, in the MAX-MST-ratio problem, given a weighted graph $G$ with positive weights, we aim to find the maximum over all bi-partitions of $V(G)$ of the combined length of the minimum spanning trees for induced subgraphs divided by the minimum spanning tree of $G$. As for the geometric setting, we introduce the MAX-EMST-ratio problem, in which given dimension $d$ and a finite point set $P$ in ${{\mathbb{R}}}^{d}$, we aim to find MAX-MST-ratio of the complete geometric graph on $P$, with Euclidean distances as weights of edges. When $d$ is a fixed number, we call the problem d-MAX-EMST-ratio (e.g. $2$-MAX-EMST-ratio, $3$-MAX-EMST-ratio). We tackle these computational problems, along with establishing bounds on the maximum and average MST-ratio, by presenting the following results:
  • The MAX-MST-ratio problem is NP-hard and even hard to approximate. (Theorems 1,2)
  • The MAX-EMST-ratio problem is NP-hard. (Theorem 3)
  • In any metric space the maximum MST-ratio is smaller than $3$. (Theorem 4)
  • There exists a ${O}(n^{2})$ time $3$-approximation algorithm for the MAX-EMST-ratio problem. (Theorem 5)
  • For every set of $n$ points in the Euclidean space, the average EMST-ratio over all colorings is at least $\frac{n-2}{n-1}$. (Theorem 7)
  • The average EMST-ratio for a set of $n$ random points uniformly distributed in $[0,1]^{d}$ tends to $d]{2}$ as $n$ goes to infinity. (Theorem 6)
1.1 Related Works
The maximum EMST-ratio is closely related to the Steiner ratio of the Euclidean space. Recall that the Steiner tree of a point set is a tree that connects them via segments of minimum total length, allowing using some extra points. The Steiner ratio, $\rho_{d}$ of ${{\mathbb{R}}}^{d}$ is the infimum over all finite point sets in ${{\mathbb{R}}}^{d}$ of the length of the Steiner tree divided by the length of the minimum spanning tree of the set. Figure 1 shows an example of three points forming an equilateral triangle of side length $1$ with Steiner ratio $\sqrt{3}/2$. Gilbert and Pollack [12] conjectured that this is the best example in the plane and therefore $\rho_{2}=\sqrt{3}/2\approx 0.866$. The best-known lower bound for $\rho_{2}$ is $0.824$, due to Chung and Graham [5]. Important to us is also a universal lower bound on $\rho_{d}$ that works for every $d$. Gilbert and Pollack [12] presented a short proof for $\rho_{d}\geq 0.5$. Later, Graham and Hwang [13] showed that for any dimension $d$ we have $\rho_{d}\geq 0.577...$, which means the Steiner tree of a point set $P$ can never be shorter than 0.577×0.577..$\times$0.577 … × —EMST(P)—.
Figure 1: The point set $P$ shown by square nodes is just the three vertices of an equilateral triangle of side length 1. The black edges form EMST(P) with length $2$. Adding an extra vertex in the center, the green edges form the minimum Steiner tree of $P$ with length $\sqrt{3}$.
Using the above bounds for $\rho_{2}$, Cultrera et al. [7] showed that the supremum over all point sets in ${{\mathbb{R}}}^{2}$ of the maximum EMST-ratio is between $2.154$ and $2.427$. Note that it is not hard to see that the infimum of the maximum EMST-ratio is $1$. Their attention then shifted to lattice point sets, where they introduced the MST-ratio (despite the infinite number of points). They subsequently demonstrated that the infimum and supremum of the maximum MST-ratios across all $2$-dimensional lattices are $1.25$ and $2$, respectively.
Dumitrescu et al. [10] proved that for any set $P$ of at least $12$ points in ${{\mathbb{R}}}^{2}$, $\gamma(P)>1$. They also showed that for $n$ points sampled uniformly at random in $[0,1]^{2}$, the maximum EMST-ratio is at least $\sqrt{2}-{\varepsilon}$, for every ${\varepsilon}>0$, with a probability that tends to $1$ as $n$ goes to infinity. A tighter analysis shows that in the $d$-dimensional unit cube, the expected value of the average EMST-ratio over all colorings tends to $d]{2}$ as $n$ goes to infinity, see Theorem 6. The proof relies on the classic result by Beardwood, Halton, and Hammersley [3], saying that the length of a minimum spanning tree of a set $P$ of $n$ points uniformly distributed in $[0,1]^{d}$ satisfies
$|EMST(P)|/n^{1-1/d}$ $\rightarrow{}\beta(d)$ (1)
with probability one, where $\beta(d)>0$ is a constant depending only on the dimension. Best known bounds for $\beta(2)$ are approximately $0.6$ as a lower bound by Avram and Bertsimas [2], and $0.707$ as an upper bound by Gilbert [11].
1.2 Outline
Section 2 focuses on the abstract graphs and presents complexity results on the MAX-MST-ratio problem. Section 3 addresses the geometric setting and proves bound for the maximum MST-ratio as well as complexity results for MAX-EMST-ratio. Section 4 explores the average EMST-ratio over all colorings of a point set. Section 5 presents the experimental results regarding the maximum, average, and bipartite EMST-ratio for random point sets. Finally, Section 6 concludes the paper by presenting specific open questions aimed at deepening the understanding of the MST-ratio.
2 Maximum MST-ratio of Abstract Graphs
In this section, we study the complexity of finding the maximum MST-ratio for abstract graphs. As the primary result of this section, we show that MAX-MST-ratio is NP-hard.
Theorem 1.
The problem MAX-MST-ratio is NP-hard.
Proof.
Let $\mathcal{I}$ be an instance of the well-known NP-complete $3$-SAT problem (see [8]), with a set of variables $X$ and set of clauses $\mathcal{C}$. Let $m=|\mathcal{C}|$ be the number of clauses of $\mathcal{I}$. Without loss of generality, we can assume that in $\mathcal{I}$ for each variable $x$ there exists a clause that contains $x$ and a clause that contains $\bar{x}$. Indeed, otherwise, one can deterministically decide the assignment of those variables and reach a smaller instance. Similarly, we can assume that in a fixed clause $C$, there is no variable $x$ that appears as $x$ and $\bar{x}$ as otherwise one can remove $C$ and again reach a smaller instance.
Using $\mathcal{I}$ we construct an instance $\mathcal{I^{\prime}}$ of MAX-MST-ratio on $14m+$ 1 vertices in the following way. For every clause $C$ we design the following clause gadget $H_{C}$:
  • We introduce a new vertex $C$.
  • For every variable $x\in X$ that belongs to $C$ we add a new vertex $x_{C}$.
  • For every variable $x\in X$ such that $\bar{x}$ is in $C$ we add a new vertex $\bar{x}_{C}$.
  • Let $S_{C}$ be the vertex set that contains $C$ and all the vertices $x_{C},\bar{x}_{C}$ for every variable $x\in X$. Note that $|S_{C}|=4$. For every pair of vertices $u,v\in S_{C}$, we set the weight of the edge $uv$ to $1$ (purple edges in Figure 2).
For every variable $x\in X$, we define the variable gadget for $x$ in this way:
  • For every $C,C^{\prime}\in\mathcal{C}$, if $C$ contains $x$ and $C^{\prime}$ contains $\bar{x}$ then we add an edge of weight $1$ from $x_{C}$ to $\bar{x}_{C^{\prime}}$ (orange edges in Figure 2).
We also add a vertex $F$ in our construction. For every clause $C$ we set the weight of the edge $FC$ to $1$. All the other edges in this graph have a weight $2$.
[…]
Figure 2: (a), (b), and (c) show the gadget for clauses $C_{1}=x\lor\bar{y}\lor z$, $C_{2}=\bar{x}\lor\bar{y}\lor z$, $C_{3}=\bar{x}\lor\bar{y}\lor\bar{z}$, respectively. (d) depicts the gadget of a variable $x\in X$. All the edges in in the figure have weight $1$.
In $\mathcal{I^{\prime}}$, given a bi-partition (i.e. a coloring) of vertices into sets $B$ and $R$, we are interested in $|MST(B)|+|MST(R)|$. Hence we refer to this as the value of this bi-partition (or coloring). We also denote the maximum possible value among all the possible colorings for $\mathcal{I^{\prime}}$ by $OPT(\mathcal{I^{\prime}})$. We remark that this is simply the product of the MST-ratio of $\mathcal{I^{\prime}}$ and the weight of the MST of $\mathcal{I^{\prime}}$.
Let $H$ be the induced subgraph of $\mathcal{I^{\prime}}$ on the edges of weight $1$. Observe that $H$ contains $m$ vertex-disjoint cliques of size four, which are simply the clause gadget $H_{C}$ for every $C\in\mathcal{C}$. We show that $\mathcal{I}$ is satisfiable iff $OPT(\mathcal{I^{\prime}})\geq 6m-2$.
Lemma 1.
If $\mathcal{I}$ is satisfiable, then $OPT(\mathcal{I^{\prime}})\geq 6m-2$.
Proof.
Let $A$ be an assignment that satisfies $\mathcal{I}$. We set the color of $F$ to blue and for every clause $C$ we set the color of the vertex $C$ to red. Now for every variable $x$:
  • if $A$ sets $x$ to true, then for every clause $C$ that contains $x$, we set the color of $x_{C}$ to blue and for every clause $C^{\prime}$ that contains $\bar{x}$, we set the color of $\bar{x}_{C^{\prime}}$ to red.
  • otherwise, for every clause $C$ that contains $x$, we set the color of $x_{C}$ to red and for every clause $C^{\prime}$ that contains $\bar{x}$, we set the color of $\bar{x}_{C^{\prime}}$ to blue.
Now we argue that the value of such coloring is $26m-$ 2. Firstly, observe that by construction, if the two endpoints of an edge of weight $1$ have the same color then they must belong to the same clause gadget. Therefore the subgraph induced by blue (and similarly red) vertices can only have edges of weight $2$ and the edges that belong to a clause gadget. Furthermore, since $A$ satisfies $\mathcal{I}$ then for every clause $C$ at least one of the vertices in $H_{C}$ is blue and by construction the vertex $C$ is colored in red. Altogether, if we consider the maximum forest on blue vertices and the maximum forest on red vertices of $H_{C}$, they have precisely two edges in total. Therefore, $MST(R)\cup MST(B)$ contains exactly $2m$ edges of weight $1$. Now since $\mathcal{I}^{\prime}$ has $14m+$ 1 vertices, $MST(R)\cup MST(B)$ have $14m-$ 1 edges in total. Therefore $MST(R)\cup MST(B)$ has exactly $(4m-1)-2m=2m-1$ edges of weight $2$ and hence we have $|MST(R)|+|MST(B)|=6m-2$ and hence the claim. ∎
Lemma 2.
If $OPT(\mathcal{I^{\prime}})\geq 6m-2$, then $\mathcal{I}$ is satisfiable.
Proof.
Let $Q$ be a coloring of $V(\mathcal{I^{\prime}})$ by red and blue such that if $R$ and $B$ are the set of blue and red vertices then $MST(R)+MST(B)\geq 6m-2$. First, assume that there is an edge $uv$ of weight $1$ such that $u$ and $v$ are colored with the same color in $Q$ but do not belong to the same clause gadget. Then $MST(R)\cup MST(B)$ contains at least $12m+$ 1 edges of weight $1$, as $H$ contains $m$ disjoint cliques of size $4$ and also the edge $uv$. Since $MST(R)\cup MST(B)$ has $14m-$ 1 edges then $|MST(R)|+|MST(B)|\leq(2m+1)+(4m-1-(2m+1))\times 2<6m-2$, a contradiction. This means that any edge of weight $1$ whose endpoints have the same color belongs to a clause gadget.
Now let us construct an assignment $A(Q)$, that satisfies $\mathcal{I}$. Without loss of generality, assume that $Q$ colors the vertex $F$ in blue. Then by the above argument for every $C\in\mathcal{C}$, the vertex $C$ has color red. Also, by the above argument for every variable $x$, the color of the vertices $x_{C}$ and $\bar{x}_{C^{\prime}}$ must be different in $Q$. Now consider a variable $x$. $A(Q)$ sets $x\leftarrow TRUE$ if and only if there exists a clause $C$ such that $Q$ colors $x_{C}$ by red or it colors $\bar{x}_{C}$ by blue. We claim that every clause in $\mathcal{I^{\prime}}$ is satisfied. Assume this is not the case and a clause $C$ is not satisfied. This implies that all the vertices in $H_{C}$ are colored in red. Now if we take any other clause $C^{\prime}$, in $MST(R)\cup MST(B)$ there are at least two edges from $H_{C^{\prime}}$. This means that $MST(R)\cup MST(B)$ contains at least $13+2(m-1)=2m+$ 1 edges of weight $1$. Thus $|MST(R)|+|MST(B)|\leq(2m+1)+(4m-1-(2m+1))\times 2<6m-2$, a contradiction. ∎
This completes the proof and hence we can conclude that computing MAX-MST-ratio is NP-hard. ∎
In Section 3, we will use the NP-hardness of the MAX-MST-ratio to demonstrate that the geometric version of the problem (i.e., the MAX-EMST-ratio) is also NP-hard. Moreover, in the general setting, the MAX-MST-ratio problem is extremely difficult to approximate. Theorem 2 presents a strong inapproximability result for computing the MAX-MST-ratio.
Theorem 2.
For every $0<\varepsilon\leq 1$, there is no polynomial time $O(n^{1-\varepsilon})$-approximation for MAX-MST-ratio, unless P=NP.
Proof.
We first prove that given an $\alpha$-approximation for MAX-MST-ratio, one can obtain a $2\alpha$-approximation for the MAX-Clique problem.
Let $G=(V,E)$ be a graph on $n$ vertices on which we wish to find the maximum clique. From this we create an instance of MAX-MST-ratio with the input graph $G^{\prime}$, where $G^{\prime}$ is a complete graph with $V(G^{\prime})=V$ and for each edge $e\in E(G^{\prime})$ its weight is $n$ if $e\in E$, and it is one otherwise.
First, we claim that given a clique $C$ of size $\ell$ in $G$, we can find a coloring of value at least $(\ell-2)n+(n-\ell)$ for $G^{\prime}$ in polynomial time. This is trivial if $\ell\leq 2$, and assuming $\ell\geq 3$ it is sufficient to assign the color red to all but one of the vertices in $C$, and blue to the remaining vertices. In this case, the sum of the weights in the MST of the red vertices is $(\ell-2)n$ and for the blue vertices the MST has exactly $(n-\ell)$ many edges and hence the claim follows. Now we show that given a coloring $Q$ of value $k$ for $G^{\prime}$, we can find a clique of size at least $\frac{\lceil k/n\rceil}{2}+1$ in $G$ in polynomial time. Let $R_{Q}$ and $B_{Q}$ be the red MST and blue MST of $G$ in the coloring $Q$. Note that given the weights in $G$, and since $R_{Q}$ and $B_{Q}$ have $n-2$ edges in total, then in total $R_{Q}$ and $B_{Q}$ have exactly $\lceil k/n\rceil$ edges of weight $n$. Without loss of generality, we can assume that $R_{Q}$ has at least $\frac{\lceil k/n\rceil}{2}$ edges of weight. Now consider $F:=(V(R_{Q}),E^{\prime\prime})$, where $E^{\prime\prime}$ is the set of edges of weight one in $R_{Q}$, and let $C_{1},\ldots,C_{r}$ be the connected components of $F$. Clearly, $r\geq\frac{\lceil k/n\rceil}{2}+1$. For every $1\leq i\leq k$, let $u_{i}$ be a vertex in $C_{i}$. As $R_{Q}$ is the MST on the set of red vertices then for every $i<j$, the weight of the edge $u_{i}u_{j}$ is $n$. Hence $\{u_{1},u_{2},\ldots,u_{k}\}$ is a clique of size $\frac{\ell}{2}+1$ in $G$.
Altogether, this shows that one can use an $\alpha$-approximation algorithm for the MAX-MST-ratio problem to find a $2\alpha$-approximation for the MAX-Clique problem. Moreover, Zuckerman [17] proved that there exists no $O(n^{1-\varepsilon})$-approximation for MAX-Clique problem unless $P=NP$. Therefore, there exists no $O(n^{1-\varepsilon})$-approximation for MAX-MST-ratio. ∎
3 Maximum MST-ratio of Geometric Point Sets
A weighted complete graph $G$ on $n$ vertices is realizable in $\mathbb{R}^{d}$ if there exist $n$ points in $\mathbb{R}^{d}$ such that the distance graph of these points is isomorphic to $G$.
We rely on a result by Dekster and Wilker [9], which shows that roughly speaking, any weighted complete graph with positive weights that do not differ too much, is realizable in Euclidean space.
Lemma 3.
[9, Theorem 2] For every positive integer $n>0$ there exists a real number $10<\lambda_{n}<$ 1 such that every complete graph $G$ on $n$ vertices whose all edges have weights in $[\lambda_{n}\ell,\ell]$, for a positive real number $\ell$ is realizable in ${{\mathbb{R}}}^{n-1}$.
We use the above lemma to increase all the weights of an instance in Theorem 1 by a fixed number in order to make it realizable in Euclidean space.
Lemma 4.
For every weighted complete graph $G$ on $n$ vertices, there exists a number $N>0$ such that if we increase the weight of every edge in $G$ by $N$, then the resulting graph is realizable in ${{\mathbb{R}}}^{n-1}$.
Proof.
Let $k=\binom{n}{2}$ and let $w_{1}\leq w_{2}\leq\ldots\leq w_{k}$ be the weights of the edges of $G$. Pick $N$ large enough such that $N>|w_{1}|$ and $N>\frac{\lambda_{n}w_{k}-w_{1}}{1-\lambda_{n}}$, where $\lambda_{n}$ is the real number taken from Lemma 3. Increasing the weights of the edges in $G$ by $N$ results in a graph $G^{\prime}$ with positive weights $N+w_{1}\leq N+w_{2}\leq\ldots\leq N+w_{k}$. $N$ being large enough guarantees $N+w_{1}\geq\lambda_{n}(N+w_{k})$. So $G^{\prime}$ satisfies the conditions of Lemma 3, and therefore is realizable in ${{\mathbb{R}}}^{n-1}$. ∎
Theorem 3.
The problem MAX-EMST-ratio is NP-hard.
Proof.
By Lemma 4, the constructed instances in Theorem 1 can be turned into a geometric graph by increasing the weights of all edges by a fixed amount. Note that this action does not change the optimal coloring of the graph in terms of the maximum MST-ratio. This is because the total number of edges in the minimum spanning trees of the two colors always falls two short of the total number of vertices. Thus, from any instance of $3$-SAT problem, we can construct an instance of the MAX-EMST-ratio problem. ∎
As established in Theorem 3, determining the maximum EMST-ratio is NP-hard. We conjecture that the problem remains NP-hard even when restricted to a specific dimension. In particular, we are interested in points in ${{\mathbb{R}}}^{2}$ and ${{\mathbb{R}}}^{3}$, as they are relevant for practical applications.
Conjecture 1.
The problems $2$-MAX-EMST-ratio and $3$-MAX-EMST are NP-hard.
Theorem 2 shows that it is even hard to approximate the maximum MST-ratio for general graphs. However, in Euclidean space, it is not difficult to approximate the maximum EMST-ratio as it cannot be too large. We show a general upper bound for the maximum MST-ratio in any metric space that leads to a simple $3$-approximation algorithm for the MAX-EMST-ratio problem. In the proof, we use the following well-known fact about the path double cover of trees.
Lemma 5 ([15]).
Given a tree $T$ with $\ell$ leaves, there exist $\ell$ (not necessarily pairwise distinct) paths in $T$ such that each edge of $T$ belongs to precisely two of these paths.
Theorem 4.
In any metric space, the maximum MST-ratio for every set of $n$ points ($n\geq 5$) is not more than $3-\frac{4}{n-1}$.
Proof.
Consider an arbitrary point set $S$ in a metric space, and let $T$ be a minimum spanning tree of $S$. To prove the statement, for every coloring $Q$ of $S$ we build spanning trees $R_{Q}$ on the set of red vertices, and $B_{Q}$ on the set of blue vertices of $Q$, such that the total weight of $R_{Q}\cup B_{Q}$ is at most $(3-\frac{4}{n-1})|T|$. Let $P^{*}$ be a path with maximum total weight among all paths in $T$ whose endpoints are not leaves of $T$. Assume, without loss of generality, that one endpoint of $P^{*}$, denoted by $r$, is red, and we hang $T$ from $r$. We traverse the tree from the root $r$ using Depth-First Search (DFS), prioritizing the vertices of $P^{*}$ when exploring branches. The starting time of a node refers to the moment it is first visited during the traversal.
For an arbitrary coloring $Q$ of $S$, we build $R_{Q}$ and $B_{Q}$ in the following way (see Figure 3):
  • We add every edge $e$ of $T$ that has both endpoints in red (in blue) to $R_{Q}$ (to $B_{Q}$).
  • Consider any red (blue) vertex $v$ in $T$ and any blue (red) child $w$ of $v$ such that there exists at least one red (blue) vertex in the descendants of $w$ (e.g. $v=r$ in Figure 3). Let $u_{1},u_{2},\ldots,u_{x}$ be the red descendants of $w$, sorted by starting time (i.e. $u_{1}$ has the earliest starting time, and $u_{x}$ the latest), for which all the internal vertices of the $wu_{i}$-path in $T$ are blue (red). We add the edges $u_{1}u_{2},u_{2}u_{3},\ldots,u_{x-1}u_{x},u_{x}v$ to $R_{Q}$ (to $B_{Q}$).
  • Let $v_{1}$, $v_{2}$, $\ldots$, $v_{y}$ be the set of blue vertices that have no blue ancestors, sorted by starting time. We add the edges $v_{1}v_{2},v_{2}v_{3},\ldots,v_{y-1}v_{y}$ to $B_{Q}$.
$r$$u_{3}$$u_{2}$$u_{1}$$w$
Figure 3: The construction of $R_{Q}$ and $B_{Q}$ (dashed edges) according to the coloring $Q$ of vertices in the proof of Theorem 4. The solid black and green edges form the tree $T$ rooted at $r$. The children of any vertex are from left to right according to the order of their starting time in DFS. The green edges are the edges of $P^{*}$.
Let $\mathcal{P}$ be the set of all $uv$-paths $P_{e}$ in $T$, where $e=uv$ is an edge in $R_{Q}\cup B_{Q}$. Observe that every edge of $T$ belongs to at most $3$ paths in $\mathcal{P}$. Namely, by construction, if $x$ is the parent of $y$ and $x$ is blue (similarly red) then the edge $xy$ appears in at most one path $P_{e_{b}}$ where $e_{b}\in B_{Q}$ and at most two paths $P_{e_{r}}$ where $e_{r}\in R_{Q}$. Furthermore, there are certain edges in $T$ that are covered by at most two paths in $\mathcal{P}$, namely the leaf edges (i.e. edges incident to a leaf) and the edges in $P^{*}$ (i.e. the green path in Figure 3).
Since the instance is metric, the weight of every edge $uv$ in $|R_{Q}|\cup|B_{Q}|$ is not more than the total weight of the corresponding $uv$-path in $\mathcal{P}$. So we have $|R_{Q}|+|B_{Q}|\leq 3|T|-W_{L}-W_{P^{*}}$, where $L$ is the set of leaf edges in $T$ and $W_{L}$ and $W_{P^{*}}$ are the total weight of the edges of $L$ and $P^{*}$, respectively. It suffices to show that $W_{P^{*}}+W_{L}\geq(\frac{4}{n-1})|T|$.
If the diameter of $T$ is less than $4$, then $P^{*}\cup L=T$, and hence the claim follows as $W_{P^{*}}+W_{L}=|T|\geq(\frac{4}{n-1})|T|$. Assume that the diameter of $T$ is at least $4$. Let $T^{\prime}$ be the tree obtained from $T$ by removing all the leaves of $T$ and let $\ell$ and $\ell^{\prime}$ be the number of leaves of $T$ and $T^{\prime}$, respectively. Note that a vertex in $T$ that is a leaf in $T^{\prime}$ must be adjacent to at least one leaf in $T$ and hence $\ell^{\prime}\leq\ell$. Furthermore as the diameter of $T$ is at least $4$, then $T^{\prime}$ has at least one non-leaf vertex which implies that $\ell+\ell^{\prime}\leq n-1$ and thus $\ell^{\prime}\leq\frac{n-1}{2}$.
Using Lemma 5, there are $\ell^{\prime}$ paths in $T^{\prime}$ such that all edges of $T^{\prime}$ belong to exactly two such paths. Hence $W_{P^{*}}$ is lower bounded by $\frac{2|T^{\prime}|}{\ell^{\prime}}=\frac{2(|T|-W_{L})}{\ell^{\prime}}\geq\frac{4(|T|-W_{L})}{n-1}$.
Therefore $W_{P^{*}}+W_{L}\geq\frac{4(|T|-W_{L})}{n-1}+W_{L}\geq\frac{4}{n-1}|T|.$
Note that the above bound is the best we can achieve in metric spaces. In fact, for every odd $n$, there exists a metric space with $n$ points (elements) for which the maximum MST-ratio is $3-\frac{4}{n-1}$. To see this, let $n=2k+1$, and define the points $v_{1},\ldots,v_{n}$. For every $i<n$, we set the weight of $v_{n}v_{i}$ to be one. For every $1\leq i\leq k$, we set the weight of $v_{2i}v_{2i-1}$ to be zero. We set the weight of any other edge to two. We first argue that the weights are metric. Note that the set of possible weights are $\{0,1,2\}$, the edges of weight one are the edges incident on $v_{n}$, and the edges of weight zero form a matching. Therefore, any triangle with edge weights $a\leq b\leq c$ has either zero or two edges of weight one and has at most one edge of weight zero. Thus $(a,b,c)$ is either $(0,1,1)$, or $(0,2,2)$, or $(1,1,2)$, or $(2,2,2)$, in any case we have $c\leq a+b$. Observe that $|MST(S)|=0\times k+1\times k=k$. If we color $v_{1},v_{3},\ldots,v_{n}$ by red all the other vertices by blue, then $|MST(R)|=k$ and $|MST(B)|=2(k-1)=2k-2$. Therefore $|MST(R)|+|MST(B)|=3k-2$ and hence the MST-ratio is $\frac{3k-2}{k}=3-\frac{2}{k}=3-\frac{4}{n-1}$.
Theorem 5.
There exists a ${O}(n^{2})$ time $3$-approximation algorithm for the MAX-EMST-ratio problem.
Proof.
By Theorem 4, the maximum EMST-ratio for an $n$-element point set is bounded from above by $3-\frac{4}{n-1}$. On the other hand, one can achieve EMST-ratio of at least $\frac{n-2}{n-1}$ by first computing the minimum spanning tree of the point set, and then coloring the vertices of the two components of it after removing the shortest edge by red and blue. The running time for computing this coloring is ${O}(n^{2})$, see [8, 16], and clearly, $\frac{n-2}{n-1}\geq\frac{1}{3}.(3-\frac{4}{n-1})$. ∎
We remark that one could achieve a weaker approximation factor $3.47$ in a simpler way, by using a trivial upper bound $\frac{2}{\rho_{d}}\leq\frac{2}{0.577...}<3.47$ on the MAX-EMST-ratio.
Indeed, if we restrict to ${{\mathbb{R}}}^{2}$ (i.e. $2$-MAX-EMST-ratio), the approximation factor in Theorem 5 can be improved to $2.427+\varepsilon$, because the upper bound for the maximum EMST-ratio in the plane is $\approx 2.427$ (see [7]). As can be observed, the coloring used to present the $3$-approximation is fairly straightforward. However, there seems to be potential for improvement by discovering more refined colorings in terms of mixedness.
We introduce the Bipartite Coloring as a good candidate to approximate the maximum EMST-ratio of a point set. By bipartite coloring, we mean the partition of a point set into two subsets labeled by colors in which every edge of the minimum spanning tree has different colors on the endpoints. Intuitively, it gives a well-distributed coloring of the point set that can be computed in ${O}(n^{2})$ time. Our experiments suggest that the bipartite coloring gives a very good approximation of the maximum EMST-ratio for random point sets in the plane, see Section 5. Although the bipartite coloring seems to be effective, it does not always produce the maximum EMST-ratio. Here is an extreme example:
Consider $12n+$ 1 points on a zigzag path in the triangular grid and call them a Triangular Chain. Figure 4 illustrates a triangular chain for $n=5$. The maximum EMST-ratio of a triangular chain with $12n+$ 1 points is $\frac{3n-2}{2n}$ while the ratio for the bipartite coloring is $\frac{2n-1}{2n}$. (By stretching the chain slightly, one can ensure a unique MST and, consequently, a unique bipartite coloring.) As $n$ increases, the proportion between the maximum EMST-ratio and the bipartite EMST-ratio approaches $1.5$.
Figure 4: An example of a point set where the maximum EMST-ratio has a significant gap from the EMST-ratio of the bipartite coloring. The left panel shows a triangular chain for $n=5$, its EMST in green, and the red and blue EMSTs in the bipartite coloring. The right panel presents the coloring that achieves the maximum EMST-ratio.
We conjecture that this is the worst approximation factor of the maximum EMST-ratio that can be obtained from bipartite coloring for a point set in the plane. Recall that if this conjecture is correct, it would lead to a $O(n^{2})$ time algorithm that achieves a $(1.5)$-approximation for the $2$-MAX-EMST-ratio problem.
Conjecture 2.
For every point set in ${{\mathbb{R}}}^{2}$, the EMST-ratio for the bipartite coloring is a $(1.5)$-approximation for the maximum EMST-ratio.
Apart from the bipartite coloring, we are interested in any improved approximation algorithm for the MAX-EXST-ratio that can be obtained by a simple coloring of a point set.
Question 1.
Is there any polynomial time algorithm for $2$-approximation of the MAX-EMST-ratio problem?
4 Average MST-ratio.
The Average EMST-ratio of a point set $P$ is the average over all $2$-colorings of $P$ of the EMST-ratio. It is automatically a lower bound for the maximum MST-ratio. Besides, it could be considered as a measure to see how close a point set is to a typical random point set, as we know the behavior of random point sets. In fact, due to Ondřej Draganov and Morteza Saghafian, the expected value for the average EMST-ratio of a random point set in a unit square is $\sqrt{2}$ in the limit. Here, we present the proof for the generalized version of this result, meaning for points within a $d$-dimensional unit cube.
Theorem 6.
For a set of $n$ random points uniformly distributed in $[0,1]^{d}$, the expected value of the average EMST-ratio tends to $d]{2}$ as $n$ goes to infinity.
Proof.
Let $n$ be large enough and $P$ be a set of $n$ random points uniformly distributed in $[0,1]^{d}$. Consider all $22^{n}-$ 2 proper bipartitions $P=B\cup R$ into blue and red points. Clearly, for $11\leq k\leq n-$ 1, in $\binom{n}{k}$ of such bipartitions $|B|=k,|R|=n-k$. Ignoring the relatively small $k$ or $n-k$, by (1), we have
$|EMST(B)|\sim\beta(d).k^{1-1/d},\hskip 14.22636pt|EMST(R)|\sim\beta(d).(n-k)^{1-1/d},\hskip 14.22636pt|EMST(P)|\sim\beta(d).n^{1-1/d}$
with probability one, where $\beta(d)$ is a constant. Thus, the average EMST-ratio over all proper bipartitions of $P$ is
$\sum_{P=B\cup R}\frac{|EMST(B)|+|EMST(R)|}{|EMST(P)|}$ $\sim\sum_{k=1}^{n-1}\frac{[\beta(d).k^{1-1/d}+\beta(d).(n-k)^{1-1/d}].\binom{n}{k}}{\beta(d).n^{1-1/d}.(2^{n}-2)}$
$=\sum_{k=1}^{n-1}\frac{[k^{1-1/d}+(n-k)^{1-1/d}].\binom{n}{k}}{n^{1-1/d}.(2^{n}-2)}$
with probability one as $n$ goes to infinity. The right-hand side is a special case of convergence of Bernstein polynomials (see [4, 14]), and therefore converges to $d]{2}$ as $n$ goes to infinity. ∎
Certainly, Theorem 4 implies that $3$ is an upper bound for the average EMST-ratio of point sets in the Euclidean space. How much can this upper bound be improved? There exist point sets in ${{\mathbb{R}}}^{2}$ with average EMST-ratio $\approx 2.154$.
[…]
Figure 5: An example of a point set with a large average EMST-ratio consists of the origin and several points clustered near the three vertices of an equilateral triangle centered at the origin. As the number of points increases, the average EMST-ratio approaches approximately $2.154$.
For instance, consider the unit circle through the origin in the plane, an equilateral triangle $abc$ on it, and then a point set $P=A^{\prime}\cup B^{\prime}\cup C^{\prime}\cup\{0\}$, where $A^{\prime}$ (resp. $B^{\prime}$, $C^{\prime}$) consists of many points that lie within distance $\varepsilon$ of $a$ (resp. $b$,$c$), for $\varepsilon$ being small enough, see Figure 5. In this case, the average EMST-ratio approaches $\frac{3+2\sqrt{3}}{3}\approx 2.154$ as the number of points increases, because with high probability, in a random two coloring, each of the sets $A^{\prime}$, $B^{\prime}$ and $C^{\prime}$ will contain both colors.
As for the lower bound, we show that the average EMST-ratio for $n$ points is always not less than $\frac{n-2}{n-1}$. This is a stronger result than the first part of Theorem 2 in [10].
Theorem 7.
For every set of $n$ points in the Euclidean space, the average EMST-ratio is at least $\frac{n-2}{n-1}$.
Proof.
Consider a set of $n$ points $P$ and let $e_{1},...,e_{n-1}$ be the edges in an Euclidean minimum spanning tree $T$ of $P$. Let us assume that the edge $e_{i}$ has weight $w_{i}$ and $w_{1}\leq w_{2}\leq\ldots\leq w_{n-1}$. Considering a coloring of $P$ in which $R$ is the set of red points and $B=P\setminus R$ is the set of blue points, we call an edge of $T$ colorful if the two endpoints of it have different colors. Let $i$ be the smallest index such that the edge $e_{i}$ is colorful. $EMST(R)\cup EMST(B)\cup\{e_{i}\}$ is a spanning tree of $P$ and therefore, its weight is at least $|EMST(P)|$. This implies $|EMST(R)|+|EMST(B)|\geq|EMST(P)|-w_{i}$.
On the other hand, considering a random coloring of $P$ with two colors, the probability for the edge $e_{i}$ being colorful is $\frac{1}{2}$, and since the $e_{i}$’s do not form a cycle, their corresponding events are independent. Thus, for a fixed index $j$, the probability that in a random coloring, the shortest colorful edge of $T$ is $e_{j}$ is $(\frac{1}{2})^{j}$. Therefore, there are precisely $2^{n-j}$ colorings of $P$ in which $j$ is the smallest index such that the edge $e_{j}$ is colorful.
By the above arguments and since we have $22^{n}-$ 2 possible (proper) colorings, the average EMST-ratio is at least:
$|MST(P)|-\sum_{i=1}^{n-1}\frac{2^{n-i}w_{i}}{2^{n}-2}.$
As $\sum_{i=1}^{n-1}2^{n-i}=2^{n}-2$ and $w_{1}\leq w_{2}\leq\ldots\leq w_{n-1}$, then:
$|MST(P)|-\sum_{i=1}^{n-1}\frac{2^{n-i}w_{i}}{2^{n}-2}\geq|MST(P)|-\frac{|MST(P)|}{n-1},$
and hence the claim. ∎
Note that Theorem 7 holds also for general graphs.
The average EMST-ratio appears to be greater than 1 in many cases, though this is not always true. The examples we found in ${{\mathbb{R}}}^{2}$ consist of point sets where $n-5\leq k\leq n-1$ points are clustered closely together, forming a core, and $1\leq n-k\leq 5$ points are sufficiently far from the core such that their pairwise distances exceed their distances to the core, see Figure 6. We believe those are the only point sets with an average EMST-ratio less than $1$.
Figure 6: A point set with average EMST-ratio smaller than $1$.
Question 2.
Can the point sets in ${{\mathbb{R}}}^{2}$ with average EMST-ratio less than $1$ be characterized?
5 Computational Experiments
This section presents the results of computational experiments comparing the average EMST-ratio, maximum EMST-ratio, and the EMST-ratio for bipartite coloring, all conducted on random point sets. The setup is as follows. In a single experiment, we consider $n$ points selected uniformly at random in the unit square, we compute the EMST-ratios for all $22^{n}-$ 2 proper colorings of this point set, and we take the maximum and the average over all colorings. Additionally, we compute the EMST-ratio for the bipartite coloring. For $n$ between $5$ and $20$, we compute the mean values of the functions (i.e. the average EMST-ratio, maximum EMST-ratio, and bipartite EMST-ratio) across 100 experiments. The resulting averages are displayed in Figure 7. According to Theorem 6, the average EMST-ratio (the red curve) tends to $\sqrt{2}$. Interestingly, the diagram in Figure 7 suggests that a similar behavior might hold for the maximum EMST-ratio (the green curve) as well, though approaching a different limit. In fact, the program did not find any instance of a random point set with a maximum MST-ratio greater than $2$ in our experiments.
Figure 7: The mean values of the maximum EMST-ratios (green), the bipartite EMST-ratios (blue), and the average EMST-ratios (red) of $100$ random sets of $n$ points, where $5\leq n\leq 20$.
Question 3.
What is the limit for the expected maximum EMST-ratio of $n$ points sampled uniformly at random in $[0,1]^{2}$, as $n$ goes to infinity?
Conjecture 3.
For $n$ random points uniformly distributed in $[0,1]^{2}$, the maximum EMST-ratio is less than $2$ with probability tending to $1$ as $n\rightarrow{\infty}$.
As for the bipartite coloring, Figure 7 clearly shows how close its EMST-ratio is to the maximum EMST-ratio in expectation. Figure 8 particularly compares the EMST-ratio for the bipartite coloring and the maximum EMST-ratio across $100000$ experiments involving $5$ to $20$ points. Moreover, the maximum EMST-ratio is never more than $1.3$ times and is rarely more than $1.1$ times the bipartite EMST-ratio across all these experiments.
Figure 8: Each point in the diagram represents a random point cloud, with the $x-$ and $y-$coordinates corresponding to the bipartite EMST-ratio and the maximum EMST-ratio, respectively.
Conjecture 4.
For $n$ random points uniformly distributed in $[0,1]^{2}$, the maximum EMST-ratio is less than $1.1$ times the bipartite EMST-ratio with probability tending to $1$ as $n$ goes to infinity.
Figure 9: In this diagram each point represents a random point cloud, with the $x-$ and $y-$coordinates corresponding to the bipartite EMST-ratio and the average EMST-ratio, respectively.
Figure 9, on the other hand, presents a comparison between the EMST-ratio for the bipartite coloring and the average EMST-ratio obtained from $100000$ experiments with point sets ranging from $5$ to $20$ points. Interestingly, in all computed experiments without exception, the bipartite EMST-ratio exceeds the average EMST-ratio. This suggests that the bipartite coloring of a point set provides a better EMST-ratio than what is expected from a random coloring. Although this is not always the case for finite point sets, we believe there are very few counterexamples.
Question 4.
Can the point sets in ${{\mathbb{R}}}^{2}$ for which the bipartite EMST-ratio is smaller than the average EMST-ratio be characterized?
6 Discussion
In addition to the conjectures and questions posed in this paper, there are several other directions worth exploring further.
Coloring with Many Crossings.
For a point set $P$ in the plane and a partition $P=R\cup B$ of $P$ into red and blue points, the Chromatic Crossing Number is defined as the number of crossings between pairs of edges, one in $EMST(R)$ and one in $EMST(B)$. Intuitively, this number can be viewed as another measure of how mixed the coloring is; the higher the chromatic crossing number, the more mingled the coloring. In fact, for any point set, it is easy to find colorings with no chromatic crossings. More interesting, however, is the Maximum Chromatic Crossing Number, which represents the highest chromatic crossing number achievable across all possible colorings of the point set.
Question 5.
How large and how small can the maximum chromatic crossing number of a set of $n$ points in the plane be?
Question 6.
Is there any relation between the chromatic crossing number and the EMST-ratio? In particular, does the coloring that maximizes the EMST-ratio tend to have a high chromatic crossing number?
Conjecture 5.
Given a point set $P$ in the plane, determining the maximum chromatic crossing number of $P$ is NP-hard.
Maximum MST Length over Subsets.
Related question to finding the maximum MST-ratio is the following. In the MAX-EMST-subset problem, given a point set $P$, we aim to find the maximum EMST length over all subsets $Q\subseteq P$. The question can be raised also in the abstract setting. Namely, in the MAX-MST-subset problem, given a weighted graph $G$, we would like to find the maximum MST length over all induced subgraphs $H\subseteq G$.
Question 7.
What can we say about the complexity of MAX-MST-subset and MAX-EMST-subset problems? Are they NP-hard?
More Colors.
Similar questions can be asked when we color a point set with $c>2$ colors. To express the MST-ratio, we consider the combined length of the minimum spanning trees of points in each color divided by the length of the minimum spanning tree of the (uncolored) point set. Theorem 4 can be generalized to obtain $12c-$ 1 as an upper bound for the MST-ratio in any metric space.
Acknowledgements. The authors thank Ondřej Draganov for his valuable insights on the topic of average EMST-ratio.

### v2

@@
-The Complexity of Maximizing the MST-ratio
+On the MST-ratio: Theoretical Bounds and Complexity of Finding the Maximum
 Abstract
-Given a finite set of red and blue points in ${{\mathbb{R}}}^{d}$, the MST-ratio is the combined length of the Euclidean minimum spanning trees of red points and of blue points divided by the length of the Euclidean minimum spanning tree of the union of them.
+Given a finite set of red and blue points in ${{\mathbb{R}}}^{d}$, the MST-ratio is defined as the total length of the Euclidean minimum spanning trees of the red points and the blue points, divided by the length of the Euclidean minimum spanning tree of their union.
 The maximum MST-ratio of a point set is the maximum MST-ratio over all non-trivial colorings of its points by red and blue.
-We prove that the problem of finding the maximum MST-ratio of a given point set is NP-hard when the dimension is a part of the input.
-Moreover, we present a ${O}(n^{2})$ running time $3$-approximation algorithm for it.
-As a part of the proof, we show that in any metric space, the maximum MST-ratio is smaller than $3$.
+We prove that finding the maximum MST-ratio of a given point set is NP-hard when the dimension is part of the input.
+Moreover, we present a quadratic-time $3$-approximation algorithm for this problem.
+As part of the proof, we show that, in any metric space, the maximum MST-ratio is smaller than $3$.
 Additionally, we study the average MST-ratio over all colorings of a set of $n$ points.
-We show that this average is always at least $\frac{n-2}{n-1}$, and for $n$ random points uniformly distributed in a $d$-dimensional unit cube, the average tends to $d]{2}$ in expectation as $n$ goes to infinity.
+We show that this average is always at least $\frac{n-2}{n-1}$, and for $n$ random points uniformly distributed in a $d$-dimensional unit cube, the average tends to $d]{2}$ in expectation as $n$ approaches infinity.
+keywords:
+Minimum Spanning Tree, Discrete and Computational Geometry, NP-hardness, Approximation Algorithms
 1 Introduction
@@
 [8] studied the interactions between color classes in a colored point set from a topological point of view.
-They introduced the concept of MST-ratio as one of the measures for mingling points with different colors in a colored point set.
-Investigating this measure opens the door to interesting discrete geometry questions related to Euclidean minimum spanning trees, or EMST for short, which is the main focus of this paper.
-For a set $P$ of $n$ points in the Euclidean space ${{\mathbb{R}}}^{d}$, the Euclidean minimum spanning tree of $P$, denoted by $EMST(P)$, is the minimum spanning tree of the complete geometric graph on $P$, where the edges have weights equal to the Euclidean distance between the two endpoints.
-For a partition $P=R\cup B$ of $P$ into red and blue points, the EMST-ratio (sometimes referred to as the MST-ratio) of $P$ and $R$ is the combined length of the Euclidean minimum spanning trees of $R$ and $B$, divided by the length of the Euclidean minimum spanning trees of $P$:
+They introduced the concept of MST-ratio as one of the measures for mingling of points with different colors in a colored point set.
+Investigating this measure opens the door to interesting discrete geometry questions related to Euclidean minimum spanning trees (EMST), which is the main focus of this paper.
+For a set $P$ of $n$ points in the Euclidean space ${{\mathbb{R}}}^{d}$, an Euclidean minimum spanning tree of $P$, denoted by $EMST(P)$, is a minimum spanning tree (MST) of the complete geometric graph on $P$ where the weight of each edge is the Euclidean distance between its endpoints.
+For a partition $P=R\cup B$ of $P$ into red and blue points, the EMST-ratio (sometimes referred to as the MST-ratio) of $P$ and $R$ is the total length of an Euclidean minimum spanning tree for each of $R$ and $B$, divided by the length of a Euclidean minimum spanning trees of $P$:
 $\mu(P,R)=\frac{|EMST(R)|+|EMST(B)|}{|EMST(P)|}.$
-How much longer can the EMSTs of two finite sets be compared to the EMST of their union?
-Given this union, $P$, we are interested in the maximum ratio, $\gamma(P)$, over all non-trivial partitions of $P$ into two sets.
+How much longer can EMSTs of two finite sets be compared to an EMST of their union?
+Given this union, $P$, we are interested in the maximum ratio, $\gamma(P)$, over all non-trivial partitioning of $P$ into two sets.
 Upper and lower bounds for $\gamma$ of certain classes of point sets can be found in [9, 13], see Section 1.1 for further details.
-However, not a great deal of research has been conducted in the area of maximum EMST-ratio for higher dimensional Euclidean space.
-Moreover, the question of whether there is an efficient algorithm to compute $\gamma$ for a given point set has yet been unanswered.
+However, limited research has been conducted on the maximum EMST-ratio in higher dimensional Euclidean spaces.
+Moreover, the question of whether there is an efficient algorithm to compute $\gamma$ for a given point set remains unanswered.
 The question can also be raised in the abstract setting.
-Namely, in the MAX-MST-ratio problem, given a weighted graph $G$ with positive weights, we aim to find the maximum over all bi-partitions of $V(G)$ of the combined length of the minimum spanning trees for induced subgraphs divided by the minimum spanning tree of $G$.
+Namely, given a weighted graph $G$ with positive weights and a bi-partition $V(G)=R\cup B$ of the vertices of $G$, we define the MST-ratio of this partition as:
+$\mu(G,R)=\frac{|MST(G[R])|+|MST(G[B])|}{|MST(G)|},$
+where $G[R]$ and $G[B]$ are the induced subgraphs of $G$ on the vertex sets $R$ and $B$, respectively, and $|H|$ is the total weight of the edges of the graph $H$.
+The maximum MST-ratio of $G$, denoted by $\gamma(G)$, is the maximum ratio over all non-trivial bi-partitions of $V(G)$.
+In the MAX-MST-ratio problem, we aim to find a partition $V(G)=R\cup B$ of the vertices of $G$ that maximizes the MST-ratio.
 As for the geometric setting, we introduce the MAX-EMST-ratio problem, in which given dimension $d$ and a finite point set $P$ in ${{\mathbb{R}}}^{d}$, we aim to find MAX-MST-ratio of the complete geometric graph on $P$, with Euclidean distances as weights of edges.
@@
 $2$-MAX-EMST-ratio, $3$-MAX-EMST-ratio).
-We tackle these computational problems, along with establishing bounds on the maximum and average MST-ratio, by presenting the following results:
-  • The MAX-MST-ratio problem is NP-hard and even hard to approximate. (Theorems 1,2)
-  • The MAX-EMST-ratio problem is NP-hard. (Theorem 3)
-  • In any metric space the maximum MST-ratio is smaller than $3$. (Theorem 4)
-  • There exists a ${O}(n^{2})$ time $3$-approximation algorithm for the MAX-EMST-ratio problem. (Theorem 5)
-  • For every set of $n$ points in the Euclidean space, the average EMST-ratio over all colorings is at least $\frac{n-2}{n-1}$. (Theorem 7)
-  • The average EMST-ratio for a set of $n$ random points uniformly distributed in $[0,1]^{d}$ tends to $d]{2}$ as $n$ goes to infinity. (Theorem 6)
+Our work establishes new bounds on the maximum and average MST-ratio.
+We also analyze the computational complexity of MAX-MST-ratio and study it through the lens of approximation algorithms.
+By an $\alpha$-approximation ($\alpha\geq 1$) for MAX-MST-ratio, we mean a polynomial time algorithm that for every weighted graph $G$ provides a bi-partition $V(G)=R\cup B$ satisfying $\alpha\cdot\mu(G,R)\geq\gamma(G)$.
+We summarize our main results as follows:
+  • The MAX-MST-ratio problem is NP-hard and hard to approximate even within a factor of $O(n^{1-\varepsilon})$. (Theorems 2.1,2.7)
+  • The MAX-EMST-ratio problem is NP-hard. (Theorem 3.4)
+  • In any metric space the maximum MST-ratio is smaller than $3$. (Theorem 3.8)
+  • There exists an ${O}(n^{2})$ time $3$-approximation algorithm for the MAX-EMST-ratio problem. (Theorem 3.10)
+  • For every set of $n$ points in Euclidean space, the average EMST-ratio over all colorings is at least $\frac{n-2}{n-1}$. (Theorem 4.3)
+  • The average EMST-ratio for a set of $n$ random points uniformly distributed in $[0,1]^{d}$ tends to $d]{2}$ as $n$ goes to infinity. (Theorem 4.1)
 1.1 Related Works
@@
 Recall that the Steiner tree of a point set is a tree that connects them via segments of minimum total length, allowing using some extra points.
-The Steiner ratio, $\rho_{d}$ of ${{\mathbb{R}}}^{d}$ is the infimum over all finite point sets in ${{\mathbb{R}}}^{d}$ of the length of the Steiner tree divided by the length of the minimum spanning tree of the set.
+The Steiner ratio, $\rho_{d}$ of ${{\mathbb{R}}}^{d}$ is the infimum over all finite point sets in ${{\mathbb{R}}}^{d}$ of the length of the Steiner tree divided by the length of a minimum spanning tree of the set.
 Figure 1 shows an example of three points forming an equilateral triangle of side length $1$ with Steiner ratio $\sqrt{3}/2$.
@@
 Gilbert and Pollack [15] presented a short proof for $\rho_{d}\geq 0.5$.
-Later, Graham and Hwang [13] showed that for any dimension $d$ we have $\rho_{d}\geq 0.577...$, which means the Steiner tree of a point set $P$ can never be shorter than 0.577×0.577..$\times$0.577 … × —EMST(P)—.
+Later, Graham and Hwang [16] showed $\rho_{d}\geq 0.577$ for any $d$.
 Figure 1: The point set $P$ shown by square nodes is just the three vertices of an equilateral triangle of side length 1.
@@
 [9] showed that the supremum over all point sets in ${{\mathbb{R}}}^{2}$ of the maximum EMST-ratio is between $2.154$ and $2.427$.
-Note that it is not hard to see that the infimum of the maximum EMST-ratio is $1$.
+It is not hard to see that the infimum of the maximum EMST-ratio is $1$.
 Their attention then shifted to lattice point sets, where they introduced the MST-ratio (despite the infinite number of points).
@@
 Dumitrescu et al.
-[10] proved that for any set $P$ of at least $12$ points in ${{\mathbb{R}}}^{2}$, $\gamma(P)>1$.
+[13] proved that for any set $P$ of at least $12$ points in ${{\mathbb{R}}}^{2}$, it holds that $\gamma(P)>1$.
 They also showed that for $n$ points sampled uniformly at random in $[0,1]^{2}$, the maximum EMST-ratio is at least $\sqrt{2}-{\varepsilon}$, for every ${\varepsilon}>0$, with a probability that tends to $1$ as $n$ goes to infinity.
-A tighter analysis shows that in the $d$-dimensional unit cube, the expected value of the average EMST-ratio over all colorings tends to $d]{2}$ as $n$ goes to infinity, see Theorem 6.
-The proof relies on the classic result by Beardwood, Halton, and Hammersley [3], saying that the length of a minimum spanning tree of a set $P$ of $n$ points uniformly distributed in $[0,1]^{d}$ satisfies
+A tighter analysis shows that in the $d$-dimensional unit cube, the expected value of the average EMST-ratio over all colorings tends to $d]{2}$ as $n$ goes to infinity, see Theorem 4.1.
+The proof relies on the classic result by Beardwood, Halton, and Hammersley [4], saying that the length of an MST of a set $P$ of $n$ points uniformly distributed in $[0,1]^{d}$ (or any bounded region of volume $1$ in ${{\mathbb{R}}}^{d}$) satisfies
 $|EMST(P)|/n^{1-1/d}\rightarrow{}\beta(d)$ (1)
-with probability one, where $\beta(d)>0$ is a constant depending only on the dimension.
-Best known bounds for $\beta(2)$ are approximately $0.6$ as a lower bound by Avram and Bertsimas [2], and $0.707$ as an upper bound by Gilbert [11].
+with probability $1$, where $\beta(d)>0$ is a constant depending only on the dimension.
+Best known lower and upper bounds for $\beta(2)$ are approximately $0.6$ by Avram and Bertsimas [2], and $0.707$ by Gilbert [14], respectively.
 1.2 Outline
 Section 2 focuses on the abstract graphs and presents complexity results on the MAX-MST-ratio problem.
-Section 3 addresses the geometric setting and proves bound for the maximum MST-ratio as well as complexity results for MAX-EMST-ratio.
-Section 4 explores the average EMST-ratio over all colorings of a point set.
+Section 3 addresses the geometric setting, establishing bounds for the maximum MST-ratio and providing complexity results for MAX-EMST-ratio.
+Section 4 explores the average EMST-ratio across all colorings of a point set.
 Section 5 presents the experimental results regarding the maximum, average, and bipartite EMST-ratio for random point sets.
-Finally, Section 6 concludes the paper by presenting specific open questions aimed at deepening the understanding of the MST-ratio.
+Finally, Section 6 concludes with open questions to deepen understanding of the MST-ratio.
 2 Maximum MST-ratio of Abstract Graphs
-In this section, we study the complexity of finding the maximum MST-ratio for abstract graphs. As the primary result of this section, we show that MAX-MST-ratio is NP-hard.
-Theorem 1.
+In this section, we study the complexity of finding the maximum MST-ratio for abstract graphs.
+As the primary result of this section, we show that MAX-MST-ratio is NP-hard.
+Throughout this section and the subsequent sections, we call an edges colorful, if the two endpoints of it have different colors, otherwise we call it monochromatic.
+Theorem 2.1.
 The problem MAX-MST-ratio is NP-hard.
-Proof.
+Proof 2.2.
 Let $\mathcal{I}$ be an instance of the well-known NP-complete $3$-SAT problem (see [10]), with a set of variables $X$ and set of clauses $\mathcal{C}$.
 Let $m=|\mathcal{C}|$ be the number of clauses of $\mathcal{I}$.
-Without loss of generality, we can assume that in $\mathcal{I}$ for each variable $x$ there exists a clause that contains $x$ and a clause that contains $\bar{x}$.
+W.l.o.g., we can assume that in $\mathcal{I}$ for each variable $x$ there exists a clause that contains $x$ and a clause that contains $\bar{x}$.
 Indeed, otherwise, one can deterministically decide the assignment of those variables and reach a smaller instance.
 Similarly, we can assume that in a fixed clause $C$, there is no variable $x$ that appears as $x$ and $\bar{x}$ as otherwise one can remove $C$ and again reach a smaller instance.
-Using $\mathcal{I}$ we construct an instance $\mathcal{I^{\prime}}$ of MAX-MST-ratio on $14m+$ 1 vertices in the following way. For every clause $C$ we design the following clause gadget $H_{C}$:
+Using $\mathcal{I}$, we construct an instance $\mathcal{I^{\prime}}$ of MAX-MST-ratio on $14m+$ 1 vertices in the following way. For every clause $C$ we design the following clause gadget $H_{C}$:
   • We introduce a new vertex $C$.
-  • For every variable $x\in X$ that belongs to $C$ we add a new vertex $x_{C}$.
-  • For every variable $x\in X$ such that $\bar{x}$ is in $C$ we add a new vertex $\bar{x}_{C}$.
+  • For every variable $x\in X$ that belongs to $C$, we add a new vertex $x_{C}$.
+  • For every variable $x\in X$ such that $\bar{x}$ is in $C$, we add a new vertex $\bar{x}_{C}$.
   • Let $S_{C}$ be the vertex set that contains $C$ and all the vertices $x_{C},\bar{x}_{C}$ for every variable $x\in X$.
 Note that $|S_{C}|=4$.
-For every pair of vertices $u,v\in S_{C}$, we set the weight of the edge $uv$ to $1$ (purple edges in Figure 2).
-For every variable $x\in X$, we define the variable gadget for $x$ in this way:
-  • For every $C,C^{\prime}\in\mathcal{C}$, if $C$ contains $x$ and $C^{\prime}$ contains $\bar{x}$ then we add an edge of weight $1$ from $x_{C}$ to $\bar{x}_{C^{\prime}}$ (orange edges in Figure 2).
-We also add a vertex $F$ in our construction. For every clause $C$ we set the weight of the edge $FC$ to $1$. All the other edges in this graph have a weight $2$.
-[…]
+For every pair of vertices $u,v\in S_{C}$, we set the weight of the edge $uv$ to $1$ (see Figure 2).
+For every variable $x\in X$, we define the variable gadget for $x$ in which for every $C,C^{\prime}\in\mathcal{C}$, if $C$ contains $x$ and $C^{\prime}$ contains $\bar{x}$, we add an edge of weight $1$ from $x_{C}$ to $\bar{x}_{C^{\prime}}$ (see Figure 2).
+We also add a vertex $F$ in our construction.
+For every clause $C$ we set the weight of the edge $FC$ to $1$.
+All the other edges in this graph have a weight $2$.
+$x_{C_{1}}$$\dots$$x_{C_{2}}$$\dots$$\dots$$x_{C_{r-1}}$$\dots$$x_{C_{r}}$$\dots$$\bar{x}_{{C^{\prime}}_{1}}$$\dots$$\bar{x}_{{C^{\prime}}_{2}}$$\dots$$\dots$$\bar{x}_{{C^{\prime}}_{t-1}}$$\dots$$\bar{x}_{{C^{\prime}}_{t}}$$\dots$(d)$C_{1}$$x_{C_{1}}$$\bar{y}_{C_{1}}$$z_{C_{1}}$(a)$C_{2}$$\bar{x}_{C_{2}}$$\bar{y}_{C_{2}}$$z_{C_{2}}$(b)$C_{3}$$\bar{x}_{C_{3}}$$\bar{y}_{C_{3}}$$\bar{z}_{C_{3}}$(c)
 Figure 2: (a), (b), and (c) show the gadget for clauses $C_{1}=x\lor\bar{y}\lor z$, $C_{2}=\bar{x}\lor\bar{y}\lor z$, $C_{3}=\bar{x}\lor\bar{y}\lor\bar{z}$, respectively.
 (d) depicts the gadget of a variable $x\in X$.
-All the edges in in the figure have weight $1$.
-In $\mathcal{I^{\prime}}$, given a bi-partition (i.e.
-a coloring) of vertices into sets $B$ and $R$, we are interested in $|MST(B)|+|MST(R)|$.
-Hence we refer to this as the value of this bi-partition (or coloring).
-We also denote the maximum possible value among all the possible colorings for $\mathcal{I^{\prime}}$ by $OPT(\mathcal{I^{\prime}})$.
-We remark that this is simply the product of the MST-ratio of $\mathcal{I^{\prime}}$ and the weight of the MST of $\mathcal{I^{\prime}}$.
+All the edges in the figure have a weight $1$.
+For $\mathcal{I^{\prime}}$, we denote the maximum possible value of $|MST(R)|+|MST(B)|$ among all the possible colorings of vertices into sets $R$ and $B$ by $OPT(\mathcal{I^{\prime}})$.
+We remark that this is simply the product of the MST-ratio of $\mathcal{I^{\prime}}$ and the weight of an MST of $\mathcal{I^{\prime}}$.
 Let $H$ be the induced subgraph of $\mathcal{I^{\prime}}$ on the edges of weight $1$.
-Observe that $H$ contains $m$ vertex-disjoint cliques of size four, which are simply the clause gadget $H_{C}$ for every $C\in\mathcal{C}$.
-We show that $\mathcal{I}$ is satisfiable iff $OPT(\mathcal{I^{\prime}})\geq 6m-2$.
-Lemma 1.
+Observe that $H$ contains $m$ vertex-disjoint cliques of size four which are simply the clause gadget $H_{C}$ for every $C\in\mathcal{C}$.
+We show that $\mathcal{I}$ is satisfiable if and only if $OPT(\mathcal{I^{\prime}})\geq 6m-2$.
+Lemma 2.3.
 If $\mathcal{I}$ is satisfiable, then $OPT(\mathcal{I^{\prime}})\geq 6m-2$.
-Proof.
+Proof 2.4.
 Let $A$ be an assignment that satisfies $\mathcal{I}$. We set the color of $F$ to blue and for every clause $C$ we set the color of the vertex $C$ to red. Now for every variable $x$:
@@
   • otherwise, for every clause $C$ that contains $x$, we set the color of $x_{C}$ to red and for every clause $C^{\prime}$ that contains $\bar{x}$, we set the color of $\bar{x}_{C^{\prime}}$ to blue.
-Now we argue that the value of such coloring is $26m-$ 2.
-Firstly, observe that by construction, if the two endpoints of an edge of weight $1$ have the same color then they must belong to the same clause gadget.
-Therefore the subgraph induced by blue (and similarly red) vertices can only have edges of weight $2$ and the edges that belong to a clause gadget.
+Now we argue that if $R$ and $B$ are respectively the set of red vertices and the set of blue vertices, then $|MST(R)|+|MST(B)|\geq 6m-2$.
+Firstly, observe that by construction, if an edge of weight $1$ is monochromatic then its endpoints must belong to the same clause gadget.
+Therefore, the subgraph induced by red (and similarly blue) vertices can only have edges of weight $2$ and edges that belong to a clause gadget.
 Furthermore, since $A$ satisfies $\mathcal{I}$ then for every clause $C$ at least one of the vertices in $H_{C}$ is blue and by construction the vertex $C$ is colored in red.
-Altogether, if we consider the maximum forest on blue vertices and the maximum forest on red vertices of $H_{C}$, they have precisely two edges in total.
-Therefore, $MST(R)\cup MST(B)$ contains exactly $2m$ edges of weight $1$.
-Now since $\mathcal{I}^{\prime}$ has $14m+$ 1 vertices, $MST(R)\cup MST(B)$ have $14m-$ 1 edges in total.
-Therefore $MST(R)\cup MST(B)$ has exactly $(4m-1)-2m=2m-1$ edges of weight $2$ and hence we have $|MST(R)|+|MST(B)|=6m-2$ and hence the claim.
-∎
-Lemma 2.
+Consequently, at most two edges in $H_{C}$ belong to $MST(R)\cup MST(B)$.
+Therefore, $MST(R)\cup MST(B)$ contains at most $2m$ edges of weight $1$.
+Since $\mathcal{I}^{\prime}$ has $14m+$ 1 vertices, $MST(R)\cup MST(B)$ has $14m-$ 1 edges.
+Thus $|MST(R)|+|MST(B)|\geq 2m+2(4m-1-2m)=6m-2$, and hence the claim.
+Lemma 2.5.
 If $OPT(\mathcal{I^{\prime}})\geq 6m-2$, then $\mathcal{I}$ is satisfiable.
-Proof.
-Let $Q$ be a coloring of $V(\mathcal{I^{\prime}})$ by red and blue such that if $R$ and $B$ are the set of blue and red vertices then $MST(R)+MST(B)\geq 6m-2$.
-First, assume that there is an edge $uv$ of weight $1$ such that $u$ and $v$ are colored with the same color in $Q$ but do not belong to the same clause gadget.
-Then $MST(R)\cup MST(B)$ contains at least $12m+$ 1 edges of weight $1$, as $H$ contains $m$ disjoint cliques of size $4$ and also the edge $uv$.
-Since $MST(R)\cup MST(B)$ has $14m-$ 1 edges then $|MST(R)|+|MST(B)|\leq(2m+1)+(4m-1-(2m+1))\times 2<6m-2$, a contradiction.
-This means that any edge of weight $1$ whose endpoints have the same color belongs to a clause gadget.
-Now let us construct an assignment $A(Q)$, that satisfies $\mathcal{I}$.
-Without loss of generality, assume that $Q$ colors the vertex $F$ in blue.
-Then by the above argument for every $C\in\mathcal{C}$, the vertex $C$ has color red.
-Also, by the above argument for every variable $x$, the color of the vertices $x_{C}$ and $\bar{x}_{C^{\prime}}$ must be different in $Q$.
+Proof 2.6.
+Let $Q$ be a coloring of $V(\mathcal{I^{\prime}})$ by red and blue such that if $R$ and $B$ are the set of red and blue vertices then $MST(R)+MST(B)\geq 6m-2$.
+The monochromatic edges of weight $1$ do not contain a forest of size $12m+$ 1.
+Otherwise, since $MST(R)\cup MST(B)$ has precisely $14m-$ 1 edges, it would have at most $22m-$ 2 edges of weight $2$, so its weight would satisfy $|MST(R)|+|MST(B)|\leq(4m-1)+2m-2=6m-3$, which is a contradiction.
+Assume that there is a monochromatic edge $uv$ of weight $1$ such that its endpoints do not belong to the same clause gadget.
+Let $E^{\prime}$ be a subset of edges that contains the edge $uv$ and exactly two monochromatic edges from each clause gadget.
+As clause gadgets are vertex-disjoint, the edges of $E^{\prime}$ induce a forest of size $12m+$ 1, a contradiction.
+This means that any monochromatic edge of weight $1$ belongs to a clause gadget.
+Now we construct an assignment $A(Q)$, that satisfies $\mathcal{I}$.
+W.l.o.g., assume that $Q$ colors the vertex $F$ in blue.
+Then by the above argument for every $C\in\mathcal{C}$, the vertex $C$ has color red, and for every variable $x$, the color of the vertices $x_{C}$ and $\bar{x}_{C^{\prime}}$ are different in $Q$.
 Now consider a variable $x$.
-$A(Q)$ sets $x\leftarrow TRUE$ if and only if there exists a clause $C$ such that $Q$ colors $x_{C}$ by red or it colors $\bar{x}_{C}$ by blue.
+$A(Q)$ sets $x$ to true if and only if there exists a clause $C$ such that $Q$ colors $x_{C}$ by blue or it colors $\bar{x}_{C}$ by red.
 We claim that every clause in $\mathcal{I^{\prime}}$ is satisfied.
 Assume this is not the case and a clause $C$ is not satisfied.
-This implies that all the vertices in $H_{C}$ are colored in red.
-Now if we take any other clause $C^{\prime}$, in $MST(R)\cup MST(B)$ there are at least two edges from $H_{C^{\prime}}$.
-This means that $MST(R)\cup MST(B)$ contains at least $13+2(m-1)=2m+$ 1 edges of weight $1$.
-Thus $|MST(R)|+|MST(B)|\leq(2m+1)+(4m-1-(2m+1))\times 2<6m-2$, a contradiction.
-∎
-This completes the proof and hence we can conclude that computing MAX-MST-ratio is NP-hard. ∎
+Then all the vertices in $H_{C}$ are colored in red.
+Let $E^{\prime}$ be a subset of edges that contains three edges from $H_{C}$ that yield a spanning tree and also for any other clause $C^{\prime}$, $E^{\prime}$ has two monochromatic edges.
+Again, $E^{\prime}$ induces a forest of size $12m+$ 1, a contradiction.
+This completes the proof. We conclude that computing MAX-MST-ratio is NP-hard.
 In Section 3, we will use the NP-hardness of the MAX-MST-ratio to demonstrate that the geometric version of the problem (i.e., the MAX-EMST-ratio) is also NP-hard.
 Moreover, in the general setting, the MAX-MST-ratio problem is extremely difficult to approximate.
-Theorem 2 presents a strong inapproximability result for computing the MAX-MST-ratio.
-Theorem 2.
-For every $0<\varepsilon\leq 1$, there is no polynomial time $O(n^{1-\varepsilon})$-approximation for MAX-MST-ratio, unless P=NP.
-Proof.
-We first prove that given an $\alpha$-approximation for MAX-MST-ratio, one can obtain a $2\alpha$-approximation for the MAX-Clique problem.
+Theorem 2.7 presents a strong inapproximability result for computing the MAX-MST-ratio.
+Theorem 2.7.
+For every $0<\varepsilon\leq 1$, there is no polynomial time $O(n^{1-\varepsilon})$-approximation for MAX-MST-ratio problem, unless P=NP.
+Proof 2.8.
+We prove that given an $\alpha$-approximation for MAX-MST-ratio, one can obtain a $2\alpha$-approximation for the MAX-Clique problem.
 Let $G=(V,E)$ be a graph on $n$ vertices on which we wish to find the maximum clique.
-From this we create an instance of MAX-MST-ratio with the input graph $G^{\prime}$, where $G^{\prime}$ is a complete graph with $V(G^{\prime})=V$ and for each edge $e\in E(G^{\prime})$ its weight is $n$ if $e\in E$, and it is one otherwise.
-First, we claim that given a clique $C$ of size $\ell$ in $G$, we can find a coloring of value at least $(\ell-2)n+(n-\ell)$ for $G^{\prime}$ in polynomial time.
-This is trivial if $\ell\leq 2$, and assuming $\ell\geq 3$ it is sufficient to assign the color red to all but one of the vertices in $C$, and blue to the remaining vertices.
-In this case, the sum of the weights in the MST of the red vertices is $(\ell-2)n$ and for the blue vertices the MST has exactly $(n-\ell)$ many edges and hence the claim follows.
-Now we show that given a coloring $Q$ of value $k$ for $G^{\prime}$, we can find a clique of size at least $\frac{\lceil k/n\rceil}{2}+1$ in $G$ in polynomial time.
-Let $R_{Q}$ and $B_{Q}$ be the red MST and blue MST of $G$ in the coloring $Q$.
-Note that given the weights in $G$, and since $R_{Q}$ and $B_{Q}$ have $n-2$ edges in total, then in total $R_{Q}$ and $B_{Q}$ have exactly $\lceil k/n\rceil$ edges of weight $n$.
-Without loss of generality, we can assume that $R_{Q}$ has at least $\frac{\lceil k/n\rceil}{2}$ edges of weight.
-Now consider $F:=(V(R_{Q}),E^{\prime\prime})$, where $E^{\prime\prime}$ is the set of edges of weight one in $R_{Q}$, and let $C_{1},\ldots,C_{r}$ be the connected components of $F$.
-Clearly, $r\geq\frac{\lceil k/n\rceil}{2}+1$.
-For every $1\leq i\leq k$, let $u_{i}$ be a vertex in $C_{i}$.
-As $R_{Q}$ is the MST on the set of red vertices then for every $i<j$, the weight of the edge $u_{i}u_{j}$ is $n$.
-Hence $\{u_{1},u_{2},\ldots,u_{k}\}$ is a clique of size $\frac{\ell}{2}+1$ in $G$.
+From this, we create an instance of MAX-MST-ratio with the input graph $G^{\prime}$, where $G^{\prime}$ is a complete graph with $V(G^{\prime})=V$ and the weight of any edge $e\in E(G^{\prime})$ is $n$ if $e\in E$, and it is $1$ otherwise.
+Given a bi-partition of vertices into sets $R$ and $B$, we refer to $|MST(R)|+|MST(B)|$ as the value of this bi-partition.
+Given a clique $C$ of size $\ell$ in $G$, we can find a coloring of value $k$ for $G^{\prime}$ in polynomial time, where $\ell\leq\lfloor\frac{k}{n}\rfloor+2$.
+This is trivial if $\ell\leq 2$, and for $\ell\geq 3$ it is sufficient to color all but one of the vertices in $C$ by red, and the remaining vertices by blue.
+In this case, the weight of any MST of the red vertices is $(\ell-2)\cdot n$ and thus $\ell\leq\lfloor\frac{k}{n}\rfloor+2$.
+Now we show that given a coloring $Q$ of value $k$ for $G^{\prime}$, we can find a clique of size at least $\frac{\lfloor k/n\rfloor}{2}+1$ in $G$ in polynomial time.
+Let $R_{Q}$ and $B_{Q}$ denote a red MST and a blue MST of $G$, respectively, in the coloring $Q$.
+Given the weights in $G^{\prime}$ belong to $\{1,n\}$, and since $R_{Q}\cup B_{Q}$ has exactly $n-2$ edges, then $R_{Q}\cup B_{Q}$ has exactly $\lfloor k/n\rfloor$ edges of weight $n$.
+W.l.o.g, we can assume that $R_{Q}$ has at least $\frac{\lfloor k/n\rfloor}{2}$ edges of weight $n$.
+Now consider $F:=(V(R_{Q}),E^{\prime\prime})$, where $E^{\prime\prime}$ is the set of edges of weight $1$ in $R_{Q}$, and let $C_{1},\ldots,C_{r}$ be the connected components of $F$.
+Clearly, $r\geq\frac{\lfloor k/n\rfloor}{2}+1$.
+For every $1\leq i\leq r$, let $u_{i}$ be a vertex in $C_{i}$.
+As $R_{Q}$ is an MST on the set of red vertices, then for every $i<j$, the weight of the edge $u_{i}u_{j}$ in $G$ is $n$.
+Hence, $\{u_{1},u_{2},\ldots,u_{r}\}$ is a clique of size $r\geq\frac{\lfloor k/n\rfloor}{2}+1$ in $G$.
 Altogether, this shows that one can use an $\alpha$-approximation algorithm for the MAX-MST-ratio problem to find a $2\alpha$-approximation for the MAX-Clique problem.
-Moreover, Zuckerman [17] proved that there exists no $O(n^{1-\varepsilon})$-approximation for MAX-Clique problem unless $P=NP$.
-Therefore, there exists no $O(n^{1-\varepsilon})$-approximation for MAX-MST-ratio.
-∎
+Zuckerman [20] proved that there exists no $O(n^{1-\varepsilon})$-approximation for MAX-Clique problem, unless $P=NP$.
+Therefore, there exists no $O(n^{1-\varepsilon})$-approximation for MAX-MST-ratio, unless $P=NP$.
 3 Maximum MST-ratio of Geometric Point Sets
+We show that MAX-EMST-ratio problem is NP-hard.
 A weighted complete graph $G$ on $n$ vertices is realizable in $\mathbb{R}^{d}$ if there exist $n$ points in $\mathbb{R}^{d}$ such that the distance graph of these points is isomorphic to $G$.
 We rely on a result by Dekster and Wilker [11], which shows that roughly speaking, any weighted complete graph with positive weights that do not differ too much, is realizable in Euclidean space.
-Lemma 3.
-[9, Theorem 2] For every positive integer $n>0$ there exists a real number $10<\lambda_{n}<$ 1 such that every complete graph $G$ on $n$ vertices whose all edges have weights in $[\lambda_{n}\ell,\ell]$, for a positive real number $\ell$ is realizable in ${{\mathbb{R}}}^{n-1}$.
-We use the above lemma to increase all the weights of an instance in Theorem 1 by a fixed number in order to make it realizable in Euclidean space.
-Lemma 4.
-For every weighted complete graph $G$ on $n$ vertices, there exists a number $N>0$ such that if we increase the weight of every edge in $G$ by $N$, then the resulting graph is realizable in ${{\mathbb{R}}}^{n-1}$.
-Proof.
+Lemma 3.1.
+[11, Theorem 2] For every positive integer $n>0$ there exists a real number $\lambda_{n}=\sqrt{1-\frac{1}{\Theta(n)}}$ such that every complete graph $G$ on $n$ vertices whose all edges have weights in $[\lambda_{n}\cdot\ell,\ell]$, for a positive real number $\ell$ is realizable in ${{\mathbb{R}}}^{n-1}$.
+We use the above lemma to increase all the weights of an instance in Theorem 2.1 by a fixed number in order to make it realizable in Euclidean space.
+Lemma 3.2.
+For every weighted complete graph $G$ on $n$ vertices, there exists a positive integer $N>0$ such that if we increase the weight of every edge in $G$ by $N$, then the resulting graph is realizable in ${{\mathbb{R}}}^{n-1}$.
+Proof 3.3.
 Let $k=\binom{n}{2}$ and let $w_{1}\leq w_{2}\leq\ldots\leq w_{k}$ be the weights of the edges of $G$.
-Pick $N$ large enough such that $N>|w_{1}|$ and $N>\frac{\lambda_{n}w_{k}-w_{1}}{1-\lambda_{n}}$, where $\lambda_{n}$ is the real number taken from Lemma 3.
+Pick $N$ such that $N>|w_{1}|$ and $N>\frac{\lambda_{n}w_{k}-w_{1}}{1-\lambda_{n}}$, where $\lambda_{n}$ is the real number in Lemma 3.1.
 Increasing the weights of the edges in $G$ by $N$ results in a graph $G^{\prime}$ with positive weights $N+w_{1}\leq N+w_{2}\leq\ldots\leq N+w_{k}$.
 $N$ being large enough guarantees $N+w_{1}\geq\lambda_{n}(N+w_{k})$.
-So $G^{\prime}$ satisfies the conditions of Lemma 3, and therefore is realizable in ${{\mathbb{R}}}^{n-1}$.
-∎
-Theorem 3.
+So $G^{\prime}$ satisfies the conditions of Lemma 3.1, and therefore is realizable in ${{\mathbb{R}}}^{n-1}$.
+Theorem 3.4.
 The problem MAX-EMST-ratio is NP-hard.
-Proof.
-By Lemma 4, the constructed instances in Theorem 1 can be turned into a geometric graph by increasing the weights of all edges by a fixed amount.
+Proof 3.5.
+By Lemma 3.2, the instances in Theorem 2.1 can be turned into a geometric graph by increasing the weights of all edges by a fixed amount.
 Note that this action does not change the optimal coloring of the graph in terms of the maximum MST-ratio.
-This is because the total number of edges in the minimum spanning trees of the two colors always falls two short of the total number of vertices.
-Thus, from any instance of $3$-SAT problem, we can construct an instance of the MAX-EMST-ratio problem.
-∎
-As established in Theorem 3, determining the maximum EMST-ratio is NP-hard.
+This is because the total number of edges in the MSTs of the two colors is always two fewer than the total number of vertices.
+Moreover, the amount we add is $O(n)$, according to Lemmas 3.1,3.2.
+Thus, from any instance of $3$-SAT problem, we can construct an instance of the MAX-EMST-ratio problem in polynomial time.
+As established in Theorem 3.4, determining the maximum EMST-ratio is NP-hard.
 We conjecture that the problem remains NP-hard even when restricted to a specific dimension.
-In particular, we are interested in points in ${{\mathbb{R}}}^{2}$ and ${{\mathbb{R}}}^{3}$, as they are relevant for practical applications.
-Conjecture 1.
-The problems $2$-MAX-EMST-ratio and $3$-MAX-EMST are NP-hard.
-Theorem 2 shows that it is even hard to approximate the maximum MST-ratio for general graphs.
+In particular, we are interested in ${{\mathbb{R}}}^{2}$ and ${{\mathbb{R}}}^{3}$, as they are relevant for practical applications.
+Conjecture 3.6.
+The problems $2$-MAX-EMST-ratio and $3$-MAX-EMST-ratio are NP-hard.
+Theorem 2.7 shows that it is even hard to approximate the maximum MST-ratio for general graphs.
 However, in Euclidean space, it is not difficult to approximate the maximum EMST-ratio as it cannot be too large.
@@
 In the proof, we use the following well-known fact about the path double cover of trees.
-Lemma 5 ([15]).
+Lemma 3.7 ([18]).
 Given a tree $T$ with $\ell$ leaves, there exist $\ell$ (not necessarily pairwise distinct) paths in $T$ such that each edge of $T$ belongs to precisely two of these paths.
-Theorem 4.
-In any metric space, the maximum MST-ratio for every set of $n$ points ($n\geq 5$) is not more than $3-\frac{4}{n-1}$.
-Proof.
-Consider an arbitrary point set $S$ in a metric space, and let $T$ be a minimum spanning tree of $S$.
-To prove the statement, for every coloring $Q$ of $S$ we build spanning trees $R_{Q}$ on the set of red vertices, and $B_{Q}$ on the set of blue vertices of $Q$, such that the total weight of $R_{Q}\cup B_{Q}$ is at most $(3-\frac{4}{n-1})|T|$.
+Theorem 3.8.
+In any metric space, the maximum MST-ratio for any set of $n\geq 5$ points is at most $3-\frac{4}{n-1}$.
+Proof 3.9.
+Consider an arbitrary point set $S$ in a metric space, and let $T$ be an MST of $S$.
 Let $P^{*}$ be a path with maximum total weight among all paths in $T$ whose endpoints are not leaves of $T$.
-Assume, without loss of generality, that one endpoint of $P^{*}$, denoted by $r$, is red, and we hang $T$ from $r$.
+Assume w.l.o.g.
+that one endpoint of $P^{*}$, denoted by $r$, is red, and we hang $T$ from $r$.
 We traverse the tree from the root $r$ using Depth-First Search (DFS), prioritizing the vertices of $P^{*}$ when exploring branches.
 The starting time of a node refers to the moment it is first visited during the traversal.
-For an arbitrary coloring $Q$ of $S$, we build $R_{Q}$ and $B_{Q}$ in the following way (see Figure 3):
+To prove the statement, for every coloring $Q$ of $S$ we build spanning trees $R_{Q}$ and $B_{Q}$ on the set of red vertices and the set of blue vertices of $Q$ respectively, such that $|R_{Q}\cup B_{Q}|$ is at most $(3-\frac{4}{n-1})|T|$.
+The way we construct $R_{Q}$ and $B_{Q}$ is as follows (see Figure 3).
   • We add every edge $e$ of $T$ that has both endpoints in red (in blue) to $R_{Q}$ (to $B_{Q}$).
@@
 $v=r$ in Figure 3).
-Let $u_{1},u_{2},\ldots,u_{x}$ be the red descendants of $w$, sorted by starting time (i.e.
-$u_{1}$ has the earliest starting time, and $u_{x}$ the latest), for which all the internal vertices of the $wu_{i}$-path in $T$ are blue (red).
+Let $u_{1},u_{2},\ldots,u_{x}$ be the red descendants of $w$ for which all the internal vertices of the $wu_{i}$-path in $T$ are blue (red).
+Assume that these vertices are sorted by starting time (i.e.
+$u_{1}$ has the earliest starting time, and $u_{x}$ has the latest).
 We add the edges $u_{1}u_{2},u_{2}u_{3},\ldots,u_{x-1}u_{x},u_{x}v$ to $R_{Q}$ (to $B_{Q}$).
@@
 $r$$u_{3}$$u_{2}$$u_{1}$$w$
-Figure 3: The construction of $R_{Q}$ and $B_{Q}$ (dashed edges) according to the coloring $Q$ of vertices in the proof of Theorem 4.
+Figure 3: The construction of $R_{Q}$ and $B_{Q}$ (dashed edges) according to the coloring $Q$ of vertices in the proof of Theorem 3.8.
 The solid black and green edges form the tree $T$ rooted at $r$.
@@
 Namely, by construction, if $x$ is the parent of $y$ and $x$ is blue (similarly red) then the edge $xy$ appears in at most one path $P_{e_{b}}$ where $e_{b}\in B_{Q}$ and at most two paths $P_{e_{r}}$ where $e_{r}\in R_{Q}$.
-Furthermore, there are certain edges in $T$ that are covered by at most two paths in $\mathcal{P}$, namely the leaf edges (i.e.
+Furthermore, certain edges in $T$ are covered by at most two paths in $\mathcal{P}$, namely the leaf edges (i.e.
 edges incident to a leaf) and the edges in $P^{*}$ (i.e.
 the green path in Figure 3).
-Since the instance is metric, the weight of every edge $uv$ in $|R_{Q}|\cup|B_{Q}|$ is not more than the total weight of the corresponding $uv$-path in $\mathcal{P}$.
+Since the instance is metric, the weight of every edge $uv$ in $R_{Q}\cup B_{Q}$ is not more than the total weight of the corresponding $uv$-path in $\mathcal{P}$.
 So we have $|R_{Q}|+|B_{Q}|\leq 3|T|-W_{L}-W_{P^{*}}$, where $L$ is the set of leaf edges in $T$ and $W_{L}$ and $W_{P^{*}}$ are the total weight of the edges of $L$ and $P^{*}$, respectively.
 It suffices to show that $W_{P^{*}}+W_{L}\geq(\frac{4}{n-1})|T|$.
-If the diameter of $T$ is less than $4$, then $P^{*}\cup L=T$, and hence the claim follows as $W_{P^{*}}+W_{L}=|T|\geq(\frac{4}{n-1})|T|$.
+If the diameter of $T$ is less than $4$, then $P^{*}\cup L=T$, and hence $W_{P^{*}}+W_{L}=|T|\geq(\frac{4}{n-1})|T|$.
 Assume that the diameter of $T$ is at least $4$.
@@
 Furthermore as the diameter of $T$ is at least $4$, then $T^{\prime}$ has at least one non-leaf vertex which implies that $\ell+\ell^{\prime}\leq n-1$ and thus $\ell^{\prime}\leq\frac{n-1}{2}$.
-Using Lemma 5, there are $\ell^{\prime}$ paths in $T^{\prime}$ such that all edges of $T^{\prime}$ belong to exactly two such paths.
+Using Lemma 3.7, there are $\ell^{\prime}$ paths in $T^{\prime}$ such that all edges of $T^{\prime}$ belong to exactly two such paths.
 Hence $W_{P^{*}}$ is lower bounded by $\frac{2|T^{\prime}|}{\ell^{\prime}}=\frac{2(|T|-W_{L})}{\ell^{\prime}}\geq\frac{4(|T|-W_{L})}{n-1}$.
@@
 Note that the above bound is the best we can achieve in metric spaces.
-In fact, for every odd $n$, there exists a metric space with $n$ points (elements) for which the maximum MST-ratio is $3-\frac{4}{n-1}$.
-To see this, let $n=2k+1$, and define the points $v_{1},\ldots,v_{n}$.
-For every $i<n$, we set the weight of $v_{n}v_{i}$ to be one.
-For every $1\leq i\leq k$, we set the weight of $v_{2i}v_{2i-1}$ to be zero.
-We set the weight of any other edge to two.
+In fact, for every odd $n$, there exist metric spaces with $n$ points (elements) for which the maximum MST-ratio is arbitrarily close to $3-\frac{4}{n-1}$.
+To see this, let $n=2k+1$, and define the point set $S=\{v_{1},\ldots,v_{n}\}$.
+For every $i<n$, we set the weight of $v_{n}v_{i}$ to be $1$.
+For every $1\leq i\leq k$, we set the weight of $v_{2i}v_{2i-1}$ to be $\varepsilon$ for a sufficiently small $\varepsilon>0$.
+Any other edge has weight $2$.
 We first argue that the weights are metric.
-Note that the set of possible weights are $\{0,1,2\}$, the edges of weight one are the edges incident on $v_{n}$, and the edges of weight zero form a matching.
-Therefore, any triangle with edge weights $a\leq b\leq c$ has either zero or two edges of weight one and has at most one edge of weight zero.
-Thus $(a,b,c)$ is either $(0,1,1)$, or $(0,2,2)$, or $(1,1,2)$, or $(2,2,2)$, in any case we have $c\leq a+b$.
-Observe that $|MST(S)|=0\times k+1\times k=k$.
-If we color $v_{1},v_{3},\ldots,v_{n}$ by red all the other vertices by blue, then $|MST(R)|=k$ and $|MST(B)|=2(k-1)=2k-2$.
-Therefore $|MST(R)|+|MST(B)|=3k-2$ and hence the MST-ratio is $\frac{3k-2}{k}=3-\frac{2}{k}=3-\frac{4}{n-1}$.
-Theorem 5.
-There exists a ${O}(n^{2})$ time $3$-approximation algorithm for the MAX-EMST-ratio problem.
-Proof.
-By Theorem 4, the maximum EMST-ratio for an $n$-element point set is bounded from above by $3-\frac{4}{n-1}$.
-On the other hand, one can achieve EMST-ratio of at least $\frac{n-2}{n-1}$ by first computing the minimum spanning tree of the point set, and then coloring the vertices of the two components of it after removing the shortest edge by red and blue.
-The running time for computing this coloring is ${O}(n^{2})$, see [8, 16], and clearly, $\frac{n-2}{n-1}\geq\frac{1}{3}.(3-\frac{4}{n-1})$.
-∎
-We remark that one could achieve a weaker approximation factor $3.47$ in a simpler way, by using a trivial upper bound $\frac{2}{\rho_{d}}\leq\frac{2}{0.577...}<3.47$ on the MAX-EMST-ratio.
+The edges of weight $1$ are the edges incident to $v_{n}$, and the edges of weight $\varepsilon$ form a matching.
+Therefore, any triangle with edge weights $a\leq b\leq c$ has either zero or two edges of weight $1$ and has at most one edge of weight $\varepsilon$.
+Thus $(a,b,c)$ is either $(\varepsilon,1,1)$, or $(\varepsilon,2,2)$, or $(1,1,2)$, or $(2,2,2)$, and in all cases we have $c\leq a+b$.
+Observe that $|MST(S)|=\varepsilon\times k+1\times k=k(1+\varepsilon)$.
+If we color $v_{1},v_{3},\ldots,v_{n}$ by red and $v_{2},v_{4},\ldots,v_{n-1}$ by blue, then $|MST(R)|=k$ and $|MST(B)|=2(k-1)=2k-2$.
+Therefore $|MST(R)|+|MST(B)|=3k-2$ and hence the MST-ratio is $\frac{3k-2}{k(1+\varepsilon)}=(3-\frac{4}{n-1})\frac{1}{1+\varepsilon}$.
+Theorem 3.10.
+There exists an ${O}(n^{2})$ time $3$-approximation algorithm for the MAX-EMST-ratio problem.
+Proof 3.11.
+By Theorem 3.8, the maximum EMST-ratio for an $n$-element point set is upper bounded by $3-\frac{4}{n-1}$.
+On the other hand, one can achieve EMST-ratio of at least $\frac{n-2}{n-1}$ by first computing a minimum spanning tree of the point set in ${O}(n^{2})$ (see [10, 19]), and then coloring the vertices of the two components of it after removing the shortest edge by red and blue.
+Finally, as $\frac{n-2}{n-1}\geq\frac{1}{3}.(3-\frac{4}{n-1})$, this is an ${O}(n^{2})$ time $3$-approximation algorithm.
+We remark that a slightly improved running time for computing an EMST is provided in [3].
+Moreover, one could achieve a weaker approximation factor $3.47$ in a simpler way using a trivial upper bound $\frac{2}{\rho_{d}}\leq\frac{2}{0.577}<3.47$ on the MAX-EMST-ratio.
 Indeed, if we restrict to ${{\mathbb{R}}}^{2}$ (i.e.
-$2$-MAX-EMST-ratio), the approximation factor in Theorem 5 can be improved to $2.427+\varepsilon$, because the upper bound for the maximum EMST-ratio in the plane is $\approx 2.427$ (see [7]).
+$2$-MAX-EMST-ratio), the approximation factor in Theorem 3.10 can be improved to $2.427$, because the upper bound for the maximum EMST-ratio in the plane is $\approx 2.427$ (see [9]), and for $n>12$ a coloring with EMST-ratio greater than one can be computed in polynomial time (see [13]).
 As can be observed, the coloring used to present the $3$-approximation is fairly straightforward.
-However, there seems to be potential for improvement by discovering more refined colorings in terms of mixedness.
-We introduce the Bipartite Coloring as a good candidate to approximate the maximum EMST-ratio of a point set.
-By bipartite coloring, we mean the partition of a point set into two subsets labeled by colors in which every edge of the minimum spanning tree has different colors on the endpoints.
-Intuitively, it gives a well-distributed coloring of the point set that can be computed in ${O}(n^{2})$ time.
-Our experiments suggest that the bipartite coloring gives a very good approximation of the maximum EMST-ratio for random point sets in the plane, see Section 5.
-Although the bipartite coloring seems to be effective, it does not always produce the maximum EMST-ratio.
-Here is an extreme example:
-Consider $12n+$ 1 points on a zigzag path in the triangular grid and call them a Triangular Chain.
-Figure 4 illustrates a triangular chain for $n=5$.
-The maximum EMST-ratio of a triangular chain with $12n+$ 1 points is $\frac{3n-2}{2n}$ while the ratio for the bipartite coloring is $\frac{2n-1}{2n}$.
-(By stretching the chain slightly, one can ensure a unique MST and, consequently, a unique bipartite coloring.) As $n$ increases, the proportion between the maximum EMST-ratio and the bipartite EMST-ratio approaches $1.5$.
-Figure 4: An example of a point set where the maximum EMST-ratio has a significant gap from the EMST-ratio of the bipartite coloring.
-The left panel shows a triangular chain for $n=5$, its EMST in green, and the red and blue EMSTs in the bipartite coloring.
+However, there seems to be room for improvement by identifying more mingled colorings.
+We introduce Bipartite Colorings as good candidates to approximate the maximum EMST-ratio of a point set.
+By a bipartite coloring of a point set, we mean considering one of the Euclidean minimum spanning trees and then partitioning the point set into two subsets labeled by colors in which every edge of the taken EMST is colorful.
+Notice that since the EMST of a point set is not necessarily unique, it may have more than one bipartite colorings, namely one for each EMST.
+Intuitively, bipartite colorings give well-distributed colorings of the point set that can be computed in ${O}(n^{2})$ time ($O(n\log n)$ in the plane, see [5]).
+Our experiments suggest that bipartite colorings give a very good approximation of the maximum EMST-ratio for random point sets in the plane (see Section 5).
+Although bipartite colorings seem to be effective, they do not always produce the maximum EMST-ratio.
+Here is an extreme example: Consider $n=2k+1$ points on a zigzag path in the triangular grid and call them a Triangular Chain.
+Figure 4 illustrates a triangular chain for $k=5$.
+The maximum EMST-ratio of a triangular chain with $12k+$ 1 points is $\frac{3k-2}{2k}$ while the ratio for a bipartite coloring is $\frac{2k-1}{2k}$.
+(By stretching the chain slightly, one can ensure a unique EMST and, consequently, a unique bipartite coloring.) As $k$ increases, the proportion between the maximum EMST-ratio and this bipartite EMST-ratio approaches $1.5$.
+Figure 4: An example of a point set where the maximum EMST-ratio has a significant gap from the EMST-ratio of a bipartite coloring.
+The left panel shows a triangular chain for $k=5$, its EMST in green, and the red and blue EMSTs in a bipartite coloring.
 The right panel presents the coloring that achieves the maximum EMST-ratio.
-We conjecture that this is the worst approximation factor of the maximum EMST-ratio that can be obtained from bipartite coloring for a point set in the plane.
-Recall that if this conjecture is correct, it would lead to a $O(n^{2})$ time algorithm that achieves a $(1.5)$-approximation for the $2$-MAX-EMST-ratio problem.
-Conjecture 2.
-For every point set in ${{\mathbb{R}}}^{2}$, the EMST-ratio for the bipartite coloring is a $(1.5)$-approximation for the maximum EMST-ratio.
-Apart from the bipartite coloring, we are interested in any improved approximation algorithm for the MAX-EXST-ratio that can be obtained by a simple coloring of a point set.
-Question 1.
-Is there any polynomial time algorithm for $2$-approximation of the MAX-EMST-ratio problem?
+We conjecture that this is the worst approximation factor of the maximum EMST-ratio that can be obtained from a bipartite coloring for a point set in the plane.
+Recall that if this conjecture is correct, it would lead to an $O(n\log n)$ time algorithm that achieves a $(1.5)$-approximation for the $2$-MAX-EMST-ratio problem.
+Conjecture 3.12.
+For every point set in ${{\mathbb{R}}}^{2}$, the EMST-ratio for any bipartite coloring is a $(1.5)$-approximation for the maximum EMST-ratio.
+Apart from the bipartite colorings, we are interested in any improved approximation algorithm for the MAX-EXST-ratio that can be obtained by a simple mixed coloring of a point set.
+Question 3.13.
+Does a polynomial-time $2$-approximation algorithm exist for MAX-EMST-ratio?
 4 Average MST-ratio.
-The Average EMST-ratio of a point set $P$ is the average over all $2$-colorings of $P$ of the EMST-ratio.
-It is automatically a lower bound for the maximum MST-ratio.
+The Average EMST-ratio of a point set $P$ is the average of EMST-ratio over all non-trivial $2$-colorings of $P$.
+It is automatically a lower bound for the maximum EMST-ratio.
 Besides, it could be considered as a measure to see how close a point set is to a typical random point set, as we know the behavior of random point sets.
-In fact, due to Ondřej Draganov and Morteza Saghafian, the expected value for the average EMST-ratio of a random point set in a unit square is $\sqrt{2}$ in the limit.
-Here, we present the proof for the generalized version of this result, meaning for points within a $d$-dimensional unit cube.
-Theorem 6.
-For a set of $n$ random points uniformly distributed in $[0,1]^{d}$, the expected value of the average EMST-ratio tends to $d]{2}$ as $n$ goes to infinity.
-Proof.
+In fact, it is proved that the expected value for the average EMST-ratio of a random point set in a unit square is $\sqrt{2}$ in the limit [12].
+We present the proof for the generalized version of this result, i.e.
+for points within a $d$-dimensional unit cube (or any bounded region in ${{\mathbb{R}}}^{d}$).
+Theorem 4.1.
+For a set of $n$ random points uniformly distributed in $[0,1]^{d}$ the expected value of the average EMST-ratio tends to $d]{2}$ as $n$ goes to infinity.
+Proof 4.2.
 Let $n$ be large enough and $P$ be a set of $n$ random points uniformly distributed in $[0,1]^{d}$.
-Consider all $22^{n}-$ 2 proper bipartitions $P=B\cup R$ into blue and red points.
-Clearly, for $11\leq k\leq n-$ 1, in $\binom{n}{k}$ of such bipartitions $|B|=k,|R|=n-k$.
-Ignoring the relatively small $k$ or $n-k$, by (1), we have
-$|EMST(B)|\sim\beta(d).k^{1-1/d},\hskip 14.22636pt|EMST(R)|\sim\beta(d).(n-k)^{1-1/d},\hskip 14.22636pt|EMST(P)|\sim\beta(d).n^{1-1/d}$
-with probability one, where $\beta(d)$ is a constant. Thus, the average EMST-ratio over all proper bipartitions of $P$ is
-$\sum_{P=B\cup R}\frac{|EMST(B)|+|EMST(R)|}{|EMST(P)|}$ $\sim\sum_{k=1}^{n-1}\frac{[\beta(d).k^{1-1/d}+\beta(d).(n-k)^{1-1/d}].\binom{n}{k}}{\beta(d).n^{1-1/d}.(2^{n}-2)}$
-$=\sum_{k=1}^{n-1}\frac{[k^{1-1/d}+(n-k)^{1-1/d}].\binom{n}{k}}{n^{1-1/d}.(2^{n}-2)}$
-with probability one as $n$ goes to infinity. The right-hand side is a special case of convergence of Bernstein polynomials (see [4, 14]), and therefore converges to $d]{2}$ as $n$ goes to infinity. ∎
-Certainly, Theorem 4 implies that $3$ is an upper bound for the average EMST-ratio of point sets in the Euclidean space.
+Consider all $22^{n}-$ 2 proper bi-partitions $P=R\cup B$ into red and blue points.
+Clearly, for $11\leq k\leq n-$ 1, in $\binom{n}{k}$ of such bi-partitions $|R|=k,|B|=n-k$.
+Ignoring the relatively small $k$ or $n-k$, by (1), we know that for every $k_{0}\leq k\leq n-k_{0}$ we have
+$|EMST(R)|\sim\beta(d).k^{1-1/d},\hskip 7.96677pt|EMST(B)|\sim\beta(d).(n-k)^{1-1/d},\hskip 7.96677pt|EMST(P)|\sim\beta(d).n^{1-1/d}$
+with probability one, where $\beta(d)$ is the constant in (1), and $k_{0}$ is an appropriate constant.
+Since $n$ is significantly greater than $k_{0}$, and the EMST-ratio is at most $3$, the contribution of the EMST-ratios for colorings with $k<k_{0}$ red points or with $n-k<k_{0}$ blue points are negligible when computing the average EMST-ratio.
+Thus, the average EMST-ratio over all proper bi-partitions of $P$ is
+$\sum_{P=R\cup B}\frac{|EMST(R)|+|EMST(B)|}{|EMST(P)|}$ $\sim\sum_{k=k_{0}}^{n-k_{0}}\frac{[\beta(d).k^{1-1/d}+\beta(d).(n-k)^{1-1/d}].\binom{n}{k}}{\beta(d).n^{1-1/d}.(2^{n}-2)}$
+$=\sum_{k=k_{0}}^{n-k_{0}}\frac{[k^{1-1/d}+(n-k)^{1-1/d}].\binom{n}{k}}{n^{1-1/d}.(2^{n}-2)}$
+with probability one as $n$ goes to infinity. The right-hand side is a special case of convergence of Bernstein polynomials (see [6, 17]), and therefore converges to $d]{2}$ as $n$ goes to infinity.
+Figure 5: A point set consists of the origin and several points clustered near the three vertices of an equilateral triangle centered at the origin.
+The corresponding minimum spanning tree is depicted.
+As the number of points increases, the average EMST-ratio approaches approximately $2.154$.
+Certainly, Theorem 3.8 implies that $3$ is an upper bound for the average EMST-ratio of point sets in Euclidean space.
 How much can this upper bound be improved?
 There exist point sets in ${{\mathbb{R}}}^{2}$ with average EMST-ratio $\approx 2.154$.
-[…]
-Figure 5: An example of a point set with a large average EMST-ratio consists of the origin and several points clustered near the three vertices of an equilateral triangle centered at the origin.
-As the number of points increases, the average EMST-ratio approaches approximately $2.154$.
-For instance, consider the unit circle through the origin in the plane, an equilateral triangle $abc$ on it, and then a point set $P=A^{\prime}\cup B^{\prime}\cup C^{\prime}\cup\{0\}$, where $A^{\prime}$ (resp.
-$B^{\prime}$, $C^{\prime}$) consists of many points that lie within distance $\varepsilon$ of $a$ (resp.
+For instance, consider the unit circle through the origin in the plane, an equilateral triangle $abc$ on it, and then a point set $P=A\cup B\cup C\cup\{0\}$, where $A$ (resp.
+$B$, $C$) consists of many points that lie within distance $\varepsilon$ of $a$ (resp.
 $b$,$c$), for $\varepsilon$ being small enough, see Figure 5.
-In this case, the average EMST-ratio approaches $\frac{3+2\sqrt{3}}{3}\approx 2.154$ as the number of points increases, because with high probability, in a random two coloring, each of the sets $A^{\prime}$, $B^{\prime}$ and $C^{\prime}$ will contain both colors.
-As for the lower bound, we show that the average EMST-ratio for $n$ points is always not less than $\frac{n-2}{n-1}$. This is a stronger result than the first part of Theorem 2 in [10].
-Theorem 7.
-For every set of $n$ points in the Euclidean space, the average EMST-ratio is at least $\frac{n-2}{n-1}$.
-Proof.
-Consider a set of $n$ points $P$ and let $e_{1},...,e_{n-1}$ be the edges in an Euclidean minimum spanning tree $T$ of $P$.
-Let us assume that the edge $e_{i}$ has weight $w_{i}$ and $w_{1}\leq w_{2}\leq\ldots\leq w_{n-1}$.
-Considering a coloring of $P$ in which $R$ is the set of red points and $B=P\setminus R$ is the set of blue points, we call an edge of $T$ colorful if the two endpoints of it have different colors.
-Let $i$ be the smallest index such that the edge $e_{i}$ is colorful.
+In this case, the average EMST-ratio approaches $\frac{3+2\sqrt{3}}{3}\approx 2.154$ as the number of points increases, because with high probability, in a random bi-partition, each of the sets $A$, $B$ and $C$ will contain both colors.
+As for the lower bound, we show that the average EMST-ratio for $n$ points is always at least $\frac{n-2}{n-1}$. This is a stronger result than the first part of Theorem 2 in [13].
+Theorem 4.3.
+For every set of $n$ points in Euclidean space, the average EMST-ratio is at least $\frac{n-2}{n-1}$.
+Proof 4.4.
+Consider a set of $n$ points $P$ and let $e_{1},...,e_{n-1}$ be the edges in a Euclidean MST, $T$, of $P$.
+Assume that the edge $e_{i}$ has weight $w_{i}$ and $w_{1}\leq w_{2}\leq\ldots\leq w_{n-1}$.
+Considering a coloring $P=R\cup B$ of $P$ into red and blue points, let $i$ be the smallest index such that the edge $e_{i}$ is colorful.
 $EMST(R)\cup EMST(B)\cup\{e_{i}\}$ is a spanning tree of $P$ and therefore, its weight is at least $|EMST(P)|$.
@@
 Therefore, there are precisely $2^{n-j}$ colorings of $P$ in which $j$ is the smallest index such that the edge $e_{j}$ is colorful.
-By the above arguments and since we have $22^{n}-$ 2 possible (proper) colorings, the average EMST-ratio is at least:
-$|MST(P)|-\sum_{i=1}^{n-1}\frac{2^{n-i}w_{i}}{2^{n}-2}.$
+By the above argument and since there are $22^{n}-$ 2 possible (proper) colorings, the average EMST-ratio is at least:
+$|EMST(P)|-\sum_{i=1}^{n-1}\frac{2^{n-i}w_{i}}{2^{n}-2}.$
 As $\sum_{i=1}^{n-1}2^{n-i}=2^{n}-2$ and $w_{1}\leq w_{2}\leq\ldots\leq w_{n-1}$, then:
-$|MST(P)|-\sum_{i=1}^{n-1}\frac{2^{n-i}w_{i}}{2^{n}-2}\geq|MST(P)|-\frac{|MST(P)|}{n-1},$
-and hence the claim. ∎
-Note that Theorem 7 holds also for general graphs.
+$|EMST(P)|-\sum_{i=1}^{n-1}\frac{2^{n-i}w_{i}}{2^{n}-2}\geq|EMST(P)|-\frac{|EMST(P)|}{n-1},$
+and hence the claim.
+Note that Theorem 4.3 holds also for general graphs.
 The average EMST-ratio appears to be greater than 1 in many cases, though this is not always true.
-The examples we found in ${{\mathbb{R}}}^{2}$ consist of point sets where $n-5\leq k\leq n-1$ points are clustered closely together, forming a core, and $1\leq n-k\leq 5$ points are sufficiently far from the core such that their pairwise distances exceed their distances to the core, see Figure 6.
-We believe those are the only point sets with an average EMST-ratio less than $1$.
+The only instances we found in ${{\mathbb{R}}}^{2}$ consist of point sets where $n-6\leq k\leq n-1$ points are clustered closely together, forming a core, and $1\leq n-k\leq 6$ points are sufficiently far from the core such that their pairwise distances are not less than their distances to the core.
+We believe those are the only point sets with an average EMST-ratio of less than $1$.
+As an example, we show that the average EMST-ratio for the point set in Figure 6 is less than $1$.
+The point set consists of $n-5$ points in a small circle through the origin whose radius is $\varepsilon\in{{o}}(\frac{1}{n2^{n}})$ (the core), along with $5$ outer points as vertices of a regular pentagon inscribed in the unit circle around the origin.
+In this example, clearly $|EMST(P)|\geq 5-{{o}}(n\varepsilon)$.
+In any bi-coloring of $P$ by red and blue if an outer point has the same color as one of the core points, then the edge connecting this outer point to a core point appears in one of the monochromatic EMSTs.
+Therefore, in most of the bi-colorings of $P$ where the core contains both colors, we have $|EMST(R)|+|EMST(B)|\leq 5+{{o}}(n\varepsilon)$.
+There are only $62$ many proper bi-colorings with a mono-chromatic core, and one can show that in each of them $|EMST(R)|+|EMST(B)|<5.1$.
+In fact, in $52$ of such colorings we even have $|EMST(R)|+|EMST(B)|<4.8$.
+Thus the average EMST-ratio for this point set is upper bounded by:
+$\frac{(2^{n}-64)\times(5+{{o}}(n\varepsilon))+52\times(4.8)+10\times(5.1)}{(2^{n}-2)\times(5-{{o}}(n\varepsilon))}<1.$
 Figure 6: A point set with average EMST-ratio smaller than $1$.
-Question 2.
-Can the point sets in ${{\mathbb{R}}}^{2}$ with average EMST-ratio less than $1$ be characterized?
+The point set consists of $n-5$ points in a very small circle through the origin (the core), and $5$ outer points as vertices of a regular pentagon inscribed in the unit circle around the origin.
+The corresponding MST is depicted.
+Question 4.5.
+Can the point sets in ${{\mathbb{R}}}^{2}$ with an average EMST-ratio $<1$ be characterized?
 5 Computational Experiments
-This section presents the results of computational experiments comparing the average EMST-ratio, maximum EMST-ratio, and the EMST-ratio for bipartite coloring, all conducted on random point sets.
+This section presents the results of computational experiments comparing the average EMST-ratio, maximum EMST-ratio, and the EMST-ratio for a bipartite coloring, all conducted on random point sets.
 The setup is as follows.
-In a single experiment, we consider $n$ points selected uniformly at random in the unit square, we compute the EMST-ratios for all $22^{n}-$ 2 proper colorings of this point set, and we take the maximum and the average over all colorings.
-Additionally, we compute the EMST-ratio for the bipartite coloring.
-For $n$ between $5$ and $20$, we compute the mean values of the functions (i.e.
-the average EMST-ratio, maximum EMST-ratio, and bipartite EMST-ratio) across 100 experiments.
+In a single experiment, we consider $n$ points selected uniformly at random in the unit square, compute the EMST-ratios for all $22^{n}-$ 2 proper colorings of this point set, and take the maximum and the average over all colorings.
+Additionally, we compute the EMST-ratio for a bipartite coloring of the sample point set, by computing an MST using Prim’s algorithm [17].
+For $n$ between $5$ and $20$, we calculate the mean values of the functions (i.e.
+the average EMST-ratio, maximum EMST-ratio, and bipartite EMST-ratio) across $500$ experiments.
 The resulting averages are displayed in Figure 7.
-According to Theorem 6, the average EMST-ratio (the red curve) tends to $\sqrt{2}$.
+According to Theorem 4.1, the average EMST-ratio (the red curve) tends to $\sqrt{2}$.
 Interestingly, the diagram in Figure 7 suggests that a similar behavior might hold for the maximum EMST-ratio (the green curve) as well, though approaching a different limit.
-In fact, the program did not find any instance of a random point set with a maximum MST-ratio greater than $2$ in our experiments.
-Figure 7: The mean values of the maximum EMST-ratios (green), the bipartite EMST-ratios (blue), and the average EMST-ratios (red) of $100$ random sets of $n$ points, where $5\leq n\leq 20$.
-Question 3.
+In fact, the program did not find any instance of a random point set with a maximum EMST-ratio greater than $2$ in our experiments.
+Figure 7: The mean values of the maximum EMST-ratios (green), the bipartite EMST-ratios (blue), and the average EMST-ratios (red) of $500$ random sets of $n$ points, where $5\leq n\leq 20$.
+Question 5.1.
 What is the limit for the expected maximum EMST-ratio of $n$ points sampled uniformly at random in $[0,1]^{2}$, as $n$ goes to infinity?
-Conjecture 3.
+Conjecture 5.2.
 For $n$ random points uniformly distributed in $[0,1]^{2}$, the maximum EMST-ratio is less than $2$ with probability tending to $1$ as $n\rightarrow{\infty}$.
 As for the bipartite coloring, Figure 7 clearly shows how close its EMST-ratio is to the maximum EMST-ratio in expectation.
-Figure 8 particularly compares the EMST-ratio for the bipartite coloring and the maximum EMST-ratio across $100000$ experiments involving $5$ to $20$ points.
-Moreover, the maximum EMST-ratio is never more than $1.3$ times and is rarely more than $1.1$ times the bipartite EMST-ratio across all these experiments.
-Figure 8: Each point in the diagram represents a random point cloud, with the $x-$ and $y-$coordinates corresponding to the bipartite EMST-ratio and the maximum EMST-ratio, respectively.
-Conjecture 4.
-For $n$ random points uniformly distributed in $[0,1]^{2}$, the maximum EMST-ratio is less than $1.1$ times the bipartite EMST-ratio with probability tending to $1$ as $n$ goes to infinity.
-Figure 9: In this diagram each point represents a random point cloud, with the $x-$ and $y-$coordinates corresponding to the bipartite EMST-ratio and the average EMST-ratio, respectively.
-Figure 9, on the other hand, presents a comparison between the EMST-ratio for the bipartite coloring and the average EMST-ratio obtained from $100000$ experiments with point sets ranging from $5$ to $20$ points.
-Interestingly, in all computed experiments without exception, the bipartite EMST-ratio exceeds the average EMST-ratio.
-This suggests that the bipartite coloring of a point set provides a better EMST-ratio than what is expected from a random coloring.
+Figure 8 particularly compares the EMST-ratio for a bipartite coloring computed using Prim’s algorithm and the maximum EMST-ratio across $10^{5}$ experiments involving $5$ to $20$ points.
+Moreover, the maximum EMST-ratio is never more than $1.3$ times and is rarely more than $1.1$ times the computed bipartite EMST-ratio across all these experiments.
+Figure 8: Each point in the diagram represents a random point cloud, with the $x$ and $y$ coordinates corresponding to a bipartite EMST-ratio of the sample point set computed using Prim’s algorithm and the maximum EMST-ratio, respectively.
+Conjecture 5.3.
+For $n$ random points uniformly distributed in $[0,1]^{2}$, the maximum EMST-ratio is less than $1.1$ times any bipartite EMST-ratio of the point set with probability tending to $1$ as $n$ goes to infinity.
+Figure 9: In this diagram each point represents a random point cloud, with the $x$ and $y$ coordinates corresponding to a bipartite EMST-ratio of the sample point set computed using Prim’s algorithm and the average EMST-ratio, respectively.
+Figure 9, on the other hand, presents a comparison between a bipartite EMST-ratio computed using Prim’s algorithm and the average EMST-ratio obtained from $10^{5}$ experiments with point sets ranging from $5$ to $20$ points.
+Interestingly, in all computed experiments without exception, the computed bipartite EMST-ratio exceeds the average EMST-ratio.
+This suggests that bipartite colorings of a point set provide better EMST-ratios than what is expected from a random coloring.
 Although this is not always the case for finite point sets, we believe there are very few counterexamples.
-Question 4.
-Can the point sets in ${{\mathbb{R}}}^{2}$ for which the bipartite EMST-ratio is smaller than the average EMST-ratio be characterized?
+Question 5.4.
+Can the point sets in ${{\mathbb{R}}}^{2}$ for which a bipartite EMST-ratio of them is smaller than the average EMST-ratio be characterized?
 6 Discussion
@@
 For a point set $P$ in the plane and a partition $P=R\cup B$ of $P$ into red and blue points, the Chromatic Crossing Number is defined as the number of crossings between pairs of edges, one in $EMST(R)$ and one in $EMST(B)$.
-Intuitively, this number can be viewed as another measure of how mixed the coloring is; the higher the chromatic crossing number, the more mingled the coloring.
-In fact, for any point set, it is easy to find colorings with no chromatic crossings.
+Intuitively, this number can be viewed as another measure of how mixed the coloring is; the higher the chromatic crossing number, the more mingled the coloring appears.
+In fact, for any point set, it is easy to find colorings with no crossings between the two MSTs.
 More interesting, however, is the Maximum Chromatic Crossing Number, which represents the highest chromatic crossing number achievable across all possible colorings of the point set.
-Question 5.
+Figure 10: An example of a point set and a bi-partition with chromatic crossing number $4$.
+Question 6.1.
 How large and how small can the maximum chromatic crossing number of a set of $n$ points in the plane be?
-Question 6.
-Is there any relation between the chromatic crossing number and the EMST-ratio? In particular, does the coloring that maximizes the EMST-ratio tend to have a high chromatic crossing number?
-Conjecture 5.
-Given a point set $P$ in the plane, determining the maximum chromatic crossing number of $P$ is NP-hard.
+Question 6.2.
+Are there relations between the chromatic crossing number and the EMST-ratio? In particular, does the coloring that maximizes the chromatic crossing number tend to have a relatively high EMST-ratio?
+Conjecture 6.3.
+Finding the maximum chromatic crossing number of a point set in the plane is NP-hard.
 Maximum MST Length over Subsets.
@@
 The question can be raised also in the abstract setting.
-Namely, in the MAX-MST-subset problem, given a weighted graph $G$, we would like to find the maximum MST length over all induced subgraphs $H\subseteq G$.
-Question 7.
-What can we say about the complexity of MAX-MST-subset and MAX-EMST-subset problems? Are they NP-hard?
+Namely, in the MAX-MST-subset problem, given a weighted graph $G$, we would like to find the maximum MST weight over all induced subgraphs $H\subseteq G$.
+Question 6.4.
+Are the problems MAX-MST-subset and MAX-EMST-subset NP-hard?
 More Colors.
-Similar questions can be asked when we color a point set with $c>2$ colors.
-To express the MST-ratio, we consider the combined length of the minimum spanning trees of points in each color divided by the length of the minimum spanning tree of the (uncolored) point set.
-Theorem 4 can be generalized to obtain $12c-$ 1 as an upper bound for the MST-ratio in any metric space.
+Similar questions can be asked for point sets with $c>2$ colors.
+To express the MST-ratio, we consider the total length of minimum spanning trees of points in each color divided by the length of a minimum spanning tree of the (uncolored) point set.
+Theorem 3.8 can be generalized to obtain $12c-$ 1 as an upper bound for the MST-ratio in any metric space.
+For points in Euclidean space, this bound can be improved for $c\geq 4$.
+In particular, as $\rho_{d}\geq 0.577$ for every $d$, the maximum EMST-ratio is upper bounded by $\frac{c}{0.577}$.
+It remains unclear how the results in [9] on lattices can be generalized to more than $2$ colors.
+Additionally, exploring the average EMST-ratio for $c$ colors is an interesting direction to pursue.
 Acknowledgements. The authors thank Ondřej Draganov for his valuable insights on the topic of average EMST-ratio.

### v3

@@
-On the MST-ratio: Theoretical Bounds and Complexity of Finding the Maximum
+5600MB Eindhoven, the Netherlands
 Abstract
 Given a finite set of red and blue points in ${{\mathbb{R}}}^{d}$, the MST-ratio is defined as the total length of the Euclidean minimum spanning trees of the red points and the blue points, divided by the length of the Euclidean minimum spanning tree of their union.
-The maximum MST-ratio of a point set is the maximum MST-ratio over all non-trivial colorings of its points by red and blue.
+The MST-ratio has recently gained attention due to its direct interpretation in topological models for studying point sets with applications in spatial biology.
+The maximum MST-ratio of a point set is the maximum MST-ratio over all proper colorings of its points by red and blue.
 We prove that finding the maximum MST-ratio of a given point set is NP-hard when the dimension is part of the input.
@@
 Keywords:
-Minimum Spanning Tree, Discrete and Computational Geometry, NP-hardness, Approximation Algorithms
+Minimum Spanning Tree NP-hardness Discrete and Computational Geometry Approximation Algorithms
 1 Introduction
@@
 [9] studied the interactions between color classes in a colored point set from a topological point of view.
-They introduced the concept of MST-ratio as one of the measures for mingling of points with different colors in a colored point set.
+To this end, they developed a framework based on the chromatic Delaunay mosaic and explored its combinatorial and topological properties; see also [4].
+Moreover, they introduced the concept of MST-ratio as one of the measures for the mingling of points with different colors in a colored point set.
 Investigating this measure opens the door to interesting discrete geometry questions related to Euclidean minimum spanning trees (EMST), which is the main focus of this paper.
-For a set $P$ of $n$ points in the Euclidean space ${{\mathbb{R}}}^{d}$, an Euclidean minimum spanning tree of $P$, denoted by $EMST(P)$, is a minimum spanning tree (MST) of the complete geometric graph on $P$ where the weight of each edge is the Euclidean distance between its endpoints.
-For a partition $P=R\cup B$ of $P$ into red and blue points, the EMST-ratio (sometimes referred to as the MST-ratio) of $P$ and $R$ is the total length of an Euclidean minimum spanning tree for each of $R$ and $B$, divided by the length of a Euclidean minimum spanning trees of $P$:
-$\mu(P,R)=\frac{|EMST(R)|+|EMST(B)|}{|EMST(P)|}.$
+For a set $P$ of $n$ points in the Euclidean space ${{\mathbb{R}}}^{d}$, an Euclidean minimum spanning tree of $P$, denoted by $\text{EMST}(P)$, is a minimum spanning tree (MST) of the complete geometric graph on $P$ where the weight of each edge is the Euclidean distance between its endpoints.
+For a partition $P=R\cup B$ of $P$ into red and blue points, the EMST-ratio (sometimes referred to as the MST-ratio) can be defined as follows.
+$\mu(P,R)=\frac{|\text{EMST}(R)|+|\text{EMST}(P\setminus R)|}{|\text{EMST}(P)|}.$
+where $|\text{EMST}(X)|$ for a point set $X$ is the length of an Euclidean minimum spanning tree of $X$.
 How much longer can EMSTs of two finite sets be compared to an EMST of their union?
-Given this union, $P$, we are interested in the maximum ratio, $\gamma(P)$, over all non-trivial partitioning of $P$ into two sets.
-Upper and lower bounds for $\gamma$ of certain classes of point sets can be found in [9, 13], see Section 1.1 for further details.
-However, limited research has been conducted on the maximum EMST-ratio in higher dimensional Euclidean spaces.
+Given a point set $P$, we are interested in the maximum ratio, $\gamma(P)$, over all proper partitionings of $P$ into two sets.
+The upper and lower bounds for $\gamma$ of certain classes of point sets can be found in [10, 14], see Section 1.1 for further details.
+However, limited research has been conducted on the maximum EMST-ratio in higher-dimensional Euclidean spaces.
 Moreover, the question of whether there is an efficient algorithm to compute $\gamma$ for a given point set remains unanswered.
 The question can also be raised in the abstract setting.
-Namely, given a weighted graph $G$ with positive weights and a bi-partition $V(G)=R\cup B$ of the vertices of $G$, we define the MST-ratio of this partition as:
-$\mu(G,R)=\frac{|MST(G[R])|+|MST(G[B])|}{|MST(G)|},$
+Namely, given a weighted complete graph $G$ with positive weights and a bi-partition $V(G)=R\cup B$ of the vertices of $G$, we define the MST-ratio of this partition as:
+$\mu(G,R)=\frac{|\text{MST}(G[R])|+|\text{MST}(G[P\setminus R])|}{|\text{MST}(G)|},$
 where $G[R]$ and $G[B]$ are the induced subgraphs of $G$ on the vertex sets $R$ and $B$, respectively, and $|H|$ is the total weight of the edges of the graph $H$.
-The maximum MST-ratio of $G$, denoted by $\gamma(G)$, is the maximum ratio over all non-trivial bi-partitions of $V(G)$.
+The maximum MST-ratio of $G$, denoted by $\gamma(G)$, is the maximum ratio over all proper bi-partitions of $V(G)$.
 In the MAX-MST-ratio problem, we aim to find a partition $V(G)=R\cup B$ of the vertices of $G$ that maximizes the MST-ratio.
-As for the geometric setting, we introduce the MAX-EMST-ratio problem, in which given dimension $d$ and a finite point set $P$ in ${{\mathbb{R}}}^{d}$, we aim to find MAX-MST-ratio of the complete geometric graph on $P$, with Euclidean distances as weights of edges.
-When $d$ is a fixed number, we call the problem d-MAX-EMST-ratio (e.g.
-$2$-MAX-EMST-ratio, $3$-MAX-EMST-ratio).
+As for the geometric setting, we introduce the MAX-EMST-ratio problem, in which given dimension $d$ and a point set $P$ in ${{\mathbb{R}}}^{d}$, we aim to find the maximum EMST-ratio of $P$.
+When $d$ is a fixed number, we call the problem d-MAX-EMST-ratio.
 Our work establishes new bounds on the maximum and average MST-ratio.
 We also analyze the computational complexity of MAX-MST-ratio and study it through the lens of approximation algorithms.
-By an $\alpha$-approximation ($\alpha\geq 1$) for MAX-MST-ratio, we mean a polynomial time algorithm that for every weighted graph $G$ provides a bi-partition $V(G)=R\cup B$ satisfying $\alpha\cdot\mu(G,R)\geq\gamma(G)$.
+By an $\alpha$-approximation ($\alpha~{}\geq~{}1$) for MAX-MST-ratio, we mean a polynomial-time algorithm that for every weighted graph $G$ provides a bi-partition $V(G)=R\cup B$ satisfying $\alpha\cdot\mu(G,R)\geq\gamma(G)$.
 We summarize our main results as follows:
-  • The MAX-MST-ratio problem is NP-hard and hard to approximate even within a factor of $O(n^{1-\varepsilon})$. (Theorems 2.1,2.7)
-  • The MAX-EMST-ratio problem is NP-hard. (Theorem 3.4)
-  • In any metric space the maximum MST-ratio is smaller than $3$. (Theorem 3.8)
-  • There exists an ${O}(n^{2})$ time $3$-approximation algorithm for the MAX-EMST-ratio problem. (Theorem 3.10)
-  • For every set of $n$ points in Euclidean space, the average EMST-ratio over all colorings is at least $\frac{n-2}{n-1}$. (Theorem 4.3)
-  • The average EMST-ratio for a set of $n$ random points uniformly distributed in $[0,1]^{d}$ tends to $d]{2}$ as $n$ goes to infinity. (Theorem 4.1)
-1.1 Related Works
+  • The MAX-MST-ratio problem is NP-hard and hard to approximate even within a factor of $O(n^{1-\varepsilon})$ (Theorems 2.1).
+  • The MAX-EMST-ratio problem is NP-hard (Theorem 3.1).
+  • In any metric space, the maximum MST-ratio is smaller than $3$ (Theorem 3.2).
+  • There exists an ${O}(n^{2})$ time $3$-approximation algorithm for the MAX-EMST-ratio problem (Theorem 3.3).
+  • For every set of $n$ points in Euclidean space, the average EMST-ratio over all colorings is at least $\frac{n-2}{n-1}$ (Theorem 4.2).
+  • The average EMST-ratio for a set of $n$ random points uniformly distributed in $[0,1]^{d}$ tends to $d]{2}$ as $n$ goes to infinity (Theorem 4.1).
+1.1 Related Work
 The maximum EMST-ratio is closely related to the Steiner ratio of the Euclidean space.
-Recall that the Steiner tree of a point set is a tree that connects them via segments of minimum total length, allowing using some extra points.
-The Steiner ratio, $\rho_{d}$ of ${{\mathbb{R}}}^{d}$ is the infimum over all finite point sets in ${{\mathbb{R}}}^{d}$ of the length of the Steiner tree divided by the length of a minimum spanning tree of the set.
-Figure 1 shows an example of three points forming an equilateral triangle of side length $1$ with Steiner ratio $\sqrt{3}/2$.
-Gilbert and Pollack [15] conjectured that this is the best example in the plane and therefore $\rho_{2}=\sqrt{3}/2\approx 0.866$.
+Recall that the Steiner tree of a point set is a tree that connects the points via segments of minimum total length, allowing using some extra points.
+The Steiner ratio, $\rho_{d}$ of ${{\mathbb{R}}}^{d}$, is the infimum over all finite point sets in ${{\mathbb{R}}}^{d}$ of the length of the Steiner tree divided by the length of an EMST of the set.
+Figure 1 shows an example of three points that form an equilateral triangle of side length $1$ with Steiner ratio $\sqrt{3}/2$.
+Gilbert and Pollack [16] conjectured that this is the most extreme example in the plane and therefore $\rho_{2}=\sqrt{3}/2\approx 0.866$.
 The best-known lower bound for $\rho_{2}$ is $0.824$, due to Chung and Graham [8].
-Important to us is also a universal lower bound on $\rho_{d}$ that works for every $d$.
+It is also important for us to have a universal lower bound on $\rho_{d}$ that holds for all $d$.
 Gilbert and Pollack [16] presented a short proof for $\rho_{d}\geq 0.5$.
 Later, Graham and Hwang [17] showed $\rho_{d}\geq 0.577$ for any $d$.
-Figure 1: The point set $P$ shown by square nodes is just the three vertices of an equilateral triangle of side length 1.
-The black edges form EMST(P) with length $2$.
+Figure 1: The point set $P$ shown by square nodes is $3$ vertices of an equilateral triangle of side length $1$.
+The black edges form $\text{EMST}(P)$ with length $2$.
 Adding an extra vertex in the center, the green edges form the minimum Steiner tree of $P$ with length $\sqrt{3}$.
@@
 The proof relies on the classic result by Beardwood, Halton, and Hammersley [5], saying that the length of an MST of a set $P$ of $n$ points uniformly distributed in $[0,1]^{d}$ (or any bounded region of volume $1$ in ${{\mathbb{R}}}^{d}$) satisfies
-$|EMST(P)|/n^{1-1/d}\rightarrow{}\beta(d)$ (1)
+$|\text{EMST}(P)|/n^{1-1/d}\rightarrow{}\beta(d)$ (1)
 with probability $1$, where $\beta(d)>0$ is a constant depending only on the dimension.
-Best known lower and upper bounds for $\beta(2)$ are approximately $0.6$ by Avram and Bertsimas [2], and $0.707$ by Gilbert [14], respectively.
+Best known lower and upper bounds for $\beta(2)$ are approximately $0.6$ by Avram and Bertsimas [2] and $0.707$ by Gilbert [15], respectively.
 1.2 Outline
-Section 2 focuses on the abstract graphs and presents complexity results on the MAX-MST-ratio problem.
-Section 3 addresses the geometric setting, establishing bounds for the maximum MST-ratio and providing complexity results for MAX-EMST-ratio.
+Section 2 focuses on abstract graphs and presents complexity results on the MAX-MST-ratio problem.
+Section 3 addresses the geometric setting, establishing bounds for the maximum MST-ratio and providing complexity results for MAX-EMST-ratio problem.
 Section 4 explores the average EMST-ratio across all colorings of a point set.
-Section 5 presents the experimental results regarding the maximum, average, and bipartite EMST-ratio for random point sets.
-Finally, Section 6 concludes with open questions to deepen understanding of the MST-ratio.
+Section 5 concludes with open questions to deepen understanding of the MST-ratio.
+Section 0.A.3 presents experimental results regarding the maximum, average, and bipartite EMST-ratio for random point sets.
 2 Maximum MST-ratio of Abstract Graphs
 In this section, we study the complexity of finding the maximum MST-ratio for abstract graphs.
-As the primary result of this section, we show that MAX-MST-ratio is NP-hard.
-Throughout this section and the subsequent sections, we call an edges colorful, if the two endpoints of it have different colors, otherwise we call it monochromatic.
-Theorem 2.1.
-The problem MAX-MST-ratio is NP-hard.
-Proof 2.2.
-Let $\mathcal{I}$ be an instance of the well-known NP-complete $3$-SAT problem (see [10]), with a set of variables $X$ and set of clauses $\mathcal{C}$.
-Let $m=|\mathcal{C}|$ be the number of clauses of $\mathcal{I}$.
-W.l.o.g., we can assume that in $\mathcal{I}$ for each variable $x$ there exists a clause that contains $x$ and a clause that contains $\bar{x}$.
-Indeed, otherwise, one can deterministically decide the assignment of those variables and reach a smaller instance.
-Similarly, we can assume that in a fixed clause $C$, there is no variable $x$ that appears as $x$ and $\bar{x}$ as otherwise one can remove $C$ and again reach a smaller instance.
-Using $\mathcal{I}$, we construct an instance $\mathcal{I^{\prime}}$ of MAX-MST-ratio on $14m+$ 1 vertices in the following way. For every clause $C$ we design the following clause gadget $H_{C}$:
-  • We introduce a new vertex $C$.
-  • For every variable $x\in X$ that belongs to $C$, we add a new vertex $x_{C}$.
-  • For every variable $x\in X$ such that $\bar{x}$ is in $C$, we add a new vertex $\bar{x}_{C}$.
-  • Let $S_{C}$ be the vertex set that contains $C$ and all the vertices $x_{C},\bar{x}_{C}$ for every variable $x\in X$.
-Note that $|S_{C}|=4$.
-For every pair of vertices $u,v\in S_{C}$, we set the weight of the edge $uv$ to $1$ (see Figure 2).
-For every variable $x\in X$, we define the variable gadget for $x$ in which for every $C,C^{\prime}\in\mathcal{C}$, if $C$ contains $x$ and $C^{\prime}$ contains $\bar{x}$, we add an edge of weight $1$ from $x_{C}$ to $\bar{x}_{C^{\prime}}$ (see Figure 2).
-We also add a vertex $F$ in our construction.
-For every clause $C$ we set the weight of the edge $FC$ to $1$.
-All the other edges in this graph have a weight $2$.
-$x_{C_{1}}$$\dots$$x_{C_{2}}$$\dots$$\dots$$x_{C_{r-1}}$$\dots$$x_{C_{r}}$$\dots$$\bar{x}_{{C^{\prime}}_{1}}$$\dots$$\bar{x}_{{C^{\prime}}_{2}}$$\dots$$\dots$$\bar{x}_{{C^{\prime}}_{t-1}}$$\dots$$\bar{x}_{{C^{\prime}}_{t}}$$\dots$(d)$C_{1}$$x_{C_{1}}$$\bar{y}_{C_{1}}$$z_{C_{1}}$(a)$C_{2}$$\bar{x}_{C_{2}}$$\bar{y}_{C_{2}}$$z_{C_{2}}$(b)$C_{3}$$\bar{x}_{C_{3}}$$\bar{y}_{C_{3}}$$\bar{z}_{C_{3}}$(c)
-Figure 2: (a), (b), and (c) show the gadget for clauses $C_{1}=x\lor\bar{y}\lor z$, $C_{2}=\bar{x}\lor\bar{y}\lor z$, $C_{3}=\bar{x}\lor\bar{y}\lor\bar{z}$, respectively.
-(d) depicts the gadget of a variable $x\in X$.
-All the edges in the figure have a weight $1$.
-For $\mathcal{I^{\prime}}$, we denote the maximum possible value of $|MST(R)|+|MST(B)|$ among all the possible colorings of vertices into sets $R$ and $B$ by $OPT(\mathcal{I^{\prime}})$.
-We remark that this is simply the product of the MST-ratio of $\mathcal{I^{\prime}}$ and the weight of an MST of $\mathcal{I^{\prime}}$.
-Let $H$ be the induced subgraph of $\mathcal{I^{\prime}}$ on the edges of weight $1$.
-Observe that $H$ contains $m$ vertex-disjoint cliques of size four which are simply the clause gadget $H_{C}$ for every $C\in\mathcal{C}$.
-We show that $\mathcal{I}$ is satisfiable if and only if $OPT(\mathcal{I^{\prime}})\geq 6m-2$.
-Lemma 2.3.
-If $\mathcal{I}$ is satisfiable, then $OPT(\mathcal{I^{\prime}})\geq 6m-2$.
-Proof 2.4.
-Let $A$ be an assignment that satisfies $\mathcal{I}$. We set the color of $F$ to blue and for every clause $C$ we set the color of the vertex $C$ to red. Now for every variable $x$:
-  • if $A$ sets $x$ to true, then for every clause $C$ that contains $x$, we set the color of $x_{C}$ to blue and for every clause $C^{\prime}$ that contains $\bar{x}$, we set the color of $\bar{x}_{C^{\prime}}$ to red.
-  • otherwise, for every clause $C$ that contains $x$, we set the color of $x_{C}$ to red and for every clause $C^{\prime}$ that contains $\bar{x}$, we set the color of $\bar{x}_{C^{\prime}}$ to blue.
-Now we argue that if $R$ and $B$ are respectively the set of red vertices and the set of blue vertices, then $|MST(R)|+|MST(B)|\geq 6m-2$.
-Firstly, observe that by construction, if an edge of weight $1$ is monochromatic then its endpoints must belong to the same clause gadget.
-Therefore, the subgraph induced by red (and similarly blue) vertices can only have edges of weight $2$ and edges that belong to a clause gadget.
-Furthermore, since $A$ satisfies $\mathcal{I}$ then for every clause $C$ at least one of the vertices in $H_{C}$ is blue and by construction the vertex $C$ is colored in red.
-Consequently, at most two edges in $H_{C}$ belong to $MST(R)\cup MST(B)$.
-Therefore, $MST(R)\cup MST(B)$ contains at most $2m$ edges of weight $1$.
-Since $\mathcal{I}^{\prime}$ has $14m+$ 1 vertices, $MST(R)\cup MST(B)$ has $14m-$ 1 edges.
-Thus $|MST(R)|+|MST(B)|\geq 2m+2(4m-1-2m)=6m-2$, and hence the claim.
-Lemma 2.5.
-If $OPT(\mathcal{I^{\prime}})\geq 6m-2$, then $\mathcal{I}$ is satisfiable.
-Proof 2.6.
-Let $Q$ be a coloring of $V(\mathcal{I^{\prime}})$ by red and blue such that if $R$ and $B$ are the set of red and blue vertices then $MST(R)+MST(B)\geq 6m-2$.
-The monochromatic edges of weight $1$ do not contain a forest of size $12m+$ 1.
-Otherwise, since $MST(R)\cup MST(B)$ has precisely $14m-$ 1 edges, it would have at most $22m-$ 2 edges of weight $2$, so its weight would satisfy $|MST(R)|+|MST(B)|\leq(4m-1)+2m-2=6m-3$, which is a contradiction.
-Assume that there is a monochromatic edge $uv$ of weight $1$ such that its endpoints do not belong to the same clause gadget.
-Let $E^{\prime}$ be a subset of edges that contains the edge $uv$ and exactly two monochromatic edges from each clause gadget.
-As clause gadgets are vertex-disjoint, the edges of $E^{\prime}$ induce a forest of size $12m+$ 1, a contradiction.
-This means that any monochromatic edge of weight $1$ belongs to a clause gadget.
-Now we construct an assignment $A(Q)$, that satisfies $\mathcal{I}$.
-W.l.o.g., assume that $Q$ colors the vertex $F$ in blue.
-Then by the above argument for every $C\in\mathcal{C}$, the vertex $C$ has color red, and for every variable $x$, the color of the vertices $x_{C}$ and $\bar{x}_{C^{\prime}}$ are different in $Q$.
-Now consider a variable $x$.
-$A(Q)$ sets $x$ to true if and only if there exists a clause $C$ such that $Q$ colors $x_{C}$ by blue or it colors $\bar{x}_{C}$ by red.
-We claim that every clause in $\mathcal{I^{\prime}}$ is satisfied.
-Assume this is not the case and a clause $C$ is not satisfied.
-Then all the vertices in $H_{C}$ are colored in red.
-Let $E^{\prime}$ be a subset of edges that contains three edges from $H_{C}$ that yield a spanning tree and also for any other clause $C^{\prime}$, $E^{\prime}$ has two monochromatic edges.
-Again, $E^{\prime}$ induces a forest of size $12m+$ 1, a contradiction.
-This completes the proof. We conclude that computing MAX-MST-ratio is NP-hard.
+Throughout this section and the subsequent sections, we call an edge colorful if the two endpoints of it have different colors; otherwise, we call it monochromatic.
+As the main result of this section, we show that not only MAX-MST-ratio is NP-hard, but it is extremely hard to approximate.
 In Section 3, we will use the NP-hardness of the MAX-MST-ratio to demonstrate that the geometric version of the problem (i.e., the MAX-EMST-ratio) is also NP-hard.
-Moreover, in the general setting, the MAX-MST-ratio problem is extremely difficult to approximate.
-Theorem 2.7 presents a strong inapproximability result for computing the MAX-MST-ratio.
-Theorem 2.7.
-For every $0<\varepsilon\leq 1$, there is no polynomial time $O(n^{1-\varepsilon})$-approximation for MAX-MST-ratio problem, unless P=NP.
-Proof 2.8.
+Theorem 2.1 presents a strong inapproximability result for the MAX-MST-ratio problem.
+Theorem 2.1
+For every $0<\varepsilon\leq 1$, there is no polynomial-time $O(n^{1-\varepsilon})$-approximation with weights restricted to $1$ and $n$ for MAX-MST-ratio problem, unless P=NP.
+Proof
 We prove that given an $\alpha$-approximation for MAX-MST-ratio, one can obtain a $2\alpha$-approximation for the MAX-Clique problem.
@@
 From this, we create an instance of MAX-MST-ratio with the input graph $G^{\prime}$, where $G^{\prime}$ is a complete graph with $V(G^{\prime})=V$ and the weight of any edge $e\in E(G^{\prime})$ is $n$ if $e\in E$, and it is $1$ otherwise.
-Given a bi-partition of vertices into sets $R$ and $B$, we refer to $|MST(R)|+|MST(B)|$ as the value of this bi-partition.
+Given a bi-partition of vertices into sets $R$ and $B$, we refer to $|\text{MST}(R)|+|\text{MST}(B)|$ as the value of this bi-partition.
 Given a clique $C$ of size $\ell$ in $G$, we can find a coloring of value $k$ for $G^{\prime}$ in polynomial time, where $\ell\leq\lfloor\frac{k}{n}\rfloor+2$.
@@
 In this case, the weight of any MST of the red vertices is $(\ell-2)\cdot n$ and thus $\ell\leq\lfloor\frac{k}{n}\rfloor+2$.
-Now we show that given a coloring $Q$ of value $k$ for $G^{\prime}$, we can find a clique of size at least $\frac{\lfloor k/n\rfloor}{2}+1$ in $G$ in polynomial time.
+Conversely, given a coloring $Q$ of value $k$ for $G^{\prime}$, we show that we can find a clique of size at least $\frac{\lfloor k/n\rfloor}{2}+1$ in $G$ in polynomial time.
 Let $R_{Q}$ and $B_{Q}$ denote a red MST and a blue MST of $G$, respectively, in the coloring $Q$.
-Given the weights in $G^{\prime}$ belong to $\{1,n\}$, and since $R_{Q}\cup B_{Q}$ has exactly $n-2$ edges, then $R_{Q}\cup B_{Q}$ has exactly $\lfloor k/n\rfloor$ edges of weight $n$.
+Given that the weights in $G^{\prime}$ belong to $\{1,n\}$ and since $R_{Q}\cup B_{Q}$ has exactly $n-2$ edges, then $R_{Q}\cup B_{Q}$ has exactly $\lfloor k/n\rfloor$ edges of weight $n$.
 W.l.o.g, we can assume that $R_{Q}$ has at least $\frac{\lfloor k/n\rfloor}{2}$ edges of weight $n$.
-Now consider $F:=(V(R_{Q}),E^{\prime\prime})$, where $E^{\prime\prime}$ is the set of edges of weight $1$ in $R_{Q}$, and let $C_{1},\ldots,C_{r}$ be the connected components of $F$.
-Clearly, $r\geq\frac{\lfloor k/n\rfloor}{2}+1$.
+Consider $F:=(V(R_{Q}),E^{\prime\prime})$, where $E^{\prime\prime}$ is the set of edges of weight $1$ in $R_{Q}$.
+Let $C_{1},\ldots,C_{r}$ be the connected components of $F$.
+Clearly, we have $r\geq\frac{\lfloor k/n\rfloor}{2}+1$.
 For every $1\leq i\leq r$, let $u_{i}$ be a vertex in $C_{i}$.
@@
 We show that MAX-EMST-ratio problem is NP-hard.
+We remark that, a weighted complete graph does not necessarily correspond to the distance graph of a point set in $\mathbb{R}^{d}$.
 A weighted complete graph $G$ on $n$ vertices is realizable in $\mathbb{R}^{d}$ if there exist $n$ points in $\mathbb{R}^{d}$ such that the distance graph of these points is isomorphic to $G$.
-We rely on a result by Dekster and Wilker [11], which shows that roughly speaking, any weighted complete graph with positive weights that do not differ too much, is realizable in Euclidean space.
-Lemma 3.1.
-[11, Theorem 2] For every positive integer $n>0$ there exists a real number $\lambda_{n}=\sqrt{1-\frac{1}{\Theta(n)}}$ such that every complete graph $G$ on $n$ vertices whose all edges have weights in $[\lambda_{n}\cdot\ell,\ell]$, for a positive real number $\ell$ is realizable in ${{\mathbb{R}}}^{n-1}$.
+We rely on a result by Dekster and Wilker [12], which shows that any weighted complete graph, whose weights belong to a small range, is realizable in Euclidean space.
+Lemma 1
+[12, Theorem 2] For every positive integer $n>0$ there exists a real number $\lambda_{n}=\sqrt{1-\frac{1}{\Theta(n)}}$ such that every complete graph $G$ on $n$ vertices whose all edges have weights in $[\lambda_{n}\cdot\ell,\ell]$, for a positive real number $\ell$ is realizable in ${{\mathbb{R}}}^{n-1}$.
 We use the above lemma to increase all the weights of an instance in Theorem 2.1 by a fixed number in order to make it realizable in Euclidean space.
-Lemma 3.2.
-For every weighted complete graph $G$ on $n$ vertices, there exists a positive integer $N>0$ such that if we increase the weight of every edge in $G$ by $N$, then the resulting graph is realizable in ${{\mathbb{R}}}^{n-1}$.
-Proof 3.3.
-Let $k=\binom{n}{2}$ and let $w_{1}\leq w_{2}\leq\ldots\leq w_{k}$ be the weights of the edges of $G$.
-Pick $N$ such that $N>|w_{1}|$ and $N>\frac{\lambda_{n}w_{k}-w_{1}}{1-\lambda_{n}}$, where $\lambda_{n}$ is the real number in Lemma 3.1.
-Increasing the weights of the edges in $G$ by $N$ results in a graph $G^{\prime}$ with positive weights $N+w_{1}\leq N+w_{2}\leq\ldots\leq N+w_{k}$.
-$N$ being large enough guarantees $N+w_{1}\geq\lambda_{n}(N+w_{k})$.
-So $G^{\prime}$ satisfies the conditions of Lemma 3.1, and therefore is realizable in ${{\mathbb{R}}}^{n-1}$.
-Theorem 3.4.
-The problem MAX-EMST-ratio is NP-hard.
-Proof 3.5.
-By Lemma 3.2, the instances in Theorem 2.1 can be turned into a geometric graph by increasing the weights of all edges by a fixed amount.
+Lemma 2
+Let $G$ be a weighted complete graph $G$ on $n$ vertices.
+Let $w_{1}$ and $w_{2}$ be the minimum and maximum weights of the edges of $G$ and let $\lambda_{n}$ be as defined in Lemma 1.
+For every $N$ such that $N\geq\max\{|w_{1}|,\frac{\lambda_{n}w_{2}-w_{1}}{1-\lambda_{n}}\}$, by increasing the weight of all edges in $G$ by $N$, we obtain a realizable graph in ${{\mathbb{R}}}^{n-1}$.
+Proof
+Increasing the weights of the edges in $G$ by $N$ results in a graph $G^{\prime}$ with positive weights in which the minimum and maximum weights are $N+w_{1}$ and $N+w_{2}$.
+As $N>\frac{\lambda_{n}w_{2}-w_{1}}{1-\lambda_{n}}$ then $N+w_{1}\geq\lambda_{n}(N+w_{2})$ and hence by Lemma 1, $G^{\prime}$ is realizable in ${{\mathbb{R}}}^{n-1}$.
+Theorem 3.1
+The MAX-EMST-ratio problem is NP-hard.
+Proof
+By Lemma 2, the instances of the MAX-MST-ratio problem with weights $1$ and $n$ can be turned into a geometric graph by increasing the weights of all edges by a fixed amount.
 Note that this action does not change the optimal coloring of the graph in terms of the maximum MST-ratio.
-This is because the total number of edges in the MSTs of the two colors is always two fewer than the total number of vertices.
-Moreover, the amount we add is $O(n)$, according to Lemmas 3.1,3.2.
-Thus, from any instance of $3$-SAT problem, we can construct an instance of the MAX-EMST-ratio problem in polynomial time.
+This is because the total number of edges in the MSTs of the two colors is always $n-2$.
+Moreover, the amount we add is $O(n^{2})$, according to Lemma 2.
+Thus, from any instance of the MAX-MST-ratio with weights $1$ and $n$, we can construct an instance of the MAX-EMST-ratio problem in polynomial time and hence by Theorem 2.1.
 As established in Theorem 3.1, determining the maximum EMST-ratio is NP-hard.
@@
 In particular, we are interested in ${{\mathbb{R}}}^{2}$ and ${{\mathbb{R}}}^{3}$, as they are relevant for practical applications.
-Conjecture 3.6.
-The problems $2$-MAX-EMST-ratio and $3$-MAX-EMST-ratio are NP-hard.
+Conjecture 1
+The $k$-MAX-EMST-ratio problem for $k\in\{2,3\}$ is NP-hard.
 Theorem 2.1 shows that it is even hard to approximate the maximum MST-ratio for general graphs.
@@
 In the proof, we use the following well-known fact about the path double cover of trees.
-Lemma 3.7 ([18]).
-Given a tree $T$ with $\ell$ leaves, there exist $\ell$ (not necessarily pairwise distinct) paths in $T$ such that each edge of $T$ belongs to precisely two of these paths.
-Theorem 3.8.
+Lemma 3 ([19])
+Given a tree $T$ with $\ell$ leaves, there exist $\ell$ (not necessarily distinct) paths in $T$ such that each edge of $T$ is in precisely two of these paths.
+Theorem 3.2
 In any metric space, the maximum MST-ratio for any set of $n\geq 5$ points is at most $3-\frac{4}{n-1}$.
-Proof 3.9.
+Proof
 Consider an arbitrary point set $S$ in a metric space, and let $T$ be an MST of $S$.
@@
 To prove the statement, for every coloring $Q$ of $S$ we build spanning trees $R_{Q}$ and $B_{Q}$ on the set of red vertices and the set of blue vertices of $Q$ respectively, such that $|R_{Q}\cup B_{Q}|$ is at most $(3-\frac{4}{n-1})|T|$.
-The way we construct $R_{Q}$ and $B_{Q}$ is as follows (see Figure 3).
-  • We add every edge $e$ of $T$ that has both endpoints in red (in blue) to $R_{Q}$ (to $B_{Q}$).
+We construct $R_{Q}$ and $B_{Q}$ in the following way (see Figure 2).
+  • We add every edge $e$ of $T$ with both endpoints in red (blue) to $R_{Q}$ ($B_{Q}$).
   • Consider any red (blue) vertex $v$ in $T$ and any blue (red) child $w$ of $v$ such that there exists at least one red (blue) vertex in the descendants of $w$ (e.g.
@@
 Note that a vertex in $T$ that is a leaf in $T^{\prime}$ must be adjacent to at least one leaf in $T$ and hence $\ell^{\prime}\leq\ell$.
-Furthermore as the diameter of $T$ is at least $4$, then $T^{\prime}$ has at least one non-leaf vertex which implies that $\ell+\ell^{\prime}\leq n-1$ and thus $\ell^{\prime}\leq\frac{n-1}{2}$.
-Using Lemma 3.7, there are $\ell^{\prime}$ paths in $T^{\prime}$ such that all edges of $T^{\prime}$ belong to exactly two such paths.
-Hence $W_{P^{*}}$ is lower bounded by $\frac{2|T^{\prime}|}{\ell^{\prime}}=\frac{2(|T|-W_{L})}{\ell^{\prime}}\geq\frac{4(|T|-W_{L})}{n-1}$.
-Therefore $W_{P^{*}}+W_{L}\geq\frac{4(|T|-W_{L})}{n-1}+W_{L}\geq\frac{4}{n-1}|T|.$
+Furthermore, as the diameter of $T$ is at least $4$, then $T^{\prime}$ has at least one non-leaf vertex which implies that $\ell+\ell^{\prime}\leq n-1$ and thus, $\ell^{\prime}\leq\frac{n-1}{2}$.
+Using Lemma 3, there are $\ell^{\prime}$ paths in $T^{\prime}$ such that all edges of $T^{\prime}$ belong to exactly two such paths.
+Hence, $W_{P^{*}}$ is lower bounded by $\frac{2|T^{\prime}|}{\ell^{\prime}}=\frac{2(|T|-W_{L})}{\ell^{\prime}}\geq\frac{4(|T|-W_{L})}{n-1}$.
+Therefore, it can be concluded that $W_{P^{*}}+W_{L}\geq\frac{4(|T|-W_{L})}{n-1}+W_{L}\geq\frac{4}{n-1}|T|.$
 Note that the above bound is the best we can achieve in metric spaces.
@@
 To see this, let $n=2k+1$, and define the point set $S=\{v_{1},\ldots,v_{n}\}$.
-For every $i<n$, we set the weight of $v_{n}v_{i}$ to be $1$.
+For every $i<n$, we set the weight of $v_{n}v_{i}$ to $1$.
 For every $1\leq i\leq k$, we set the weight of $v_{2i}v_{2i-1}$ to be $\varepsilon$ for a sufficiently small $\varepsilon>0$.
@@
 The edges of weight $1$ are the edges incident to $v_{n}$, and the edges of weight $\varepsilon$ form a matching.
-Therefore, any triangle with edge weights $a\leq b\leq c$ has either zero or two edges of weight $1$ and has at most one edge of weight $\varepsilon$.
-Thus $(a,b,c)$ is either $(\varepsilon,1,1)$, or $(\varepsilon,2,2)$, or $(1,1,2)$, or $(2,2,2)$, and in all cases we have $c\leq a+b$.
-Observe that $|MST(S)|=\varepsilon\times k+1\times k=k(1+\varepsilon)$.
-If we color $v_{1},v_{3},\ldots,v_{n}$ by red and $v_{2},v_{4},\ldots,v_{n-1}$ by blue, then $|MST(R)|=k$ and $|MST(B)|=2(k-1)=2k-2$.
-Therefore $|MST(R)|+|MST(B)|=3k-2$ and hence the MST-ratio is $\frac{3k-2}{k(1+\varepsilon)}=(3-\frac{4}{n-1})\frac{1}{1+\varepsilon}$.
-Theorem 3.10.
-There exists an ${O}(n^{2})$ time $3$-approximation algorithm for the MAX-EMST-ratio problem.
-Proof 3.11.
+Therefore, any triangle with edge weights $a\leq b\leq c$ has zero or two edges of weight $1$ and has at most one edge of weight $\varepsilon$.
+Thus, $(a,b,c)$ is in $\{(\varepsilon,1,1),(\varepsilon,2,2),(1,1,2),(2,2,2)\}$, and in all cases we have $c\leq a+b$.
+Observe that $|\text{MST}(S)|=\varepsilon\times k+1\times k=k(1+\varepsilon)$.
+If we color $v_{1},v_{3},\ldots,v_{n}$ by red and $v_{2},v_{4},\ldots,v_{n-1}$ by blue, then $|\text{MST}(R)|=k$ and $|\text{MST}(B)|=2(k-1)=2k-2$.
+Thus, $|\text{MST}(R)|+|\text{MST}(B)|=3k-2$ and hence the MST-ratio is $\frac{3k-2}{k(1+\varepsilon)}=(3-\frac{4}{n-1})\frac{1}{1+\varepsilon}$.
+Theorem 3.3
+There exists an ${O}(n^{2})$-time $3$-approximation algorithm for the MAX-EMST-ratio problem.
+Proof
 By Theorem 3.2, the maximum EMST-ratio for an $n$-element point set is upper bounded by $3-\frac{4}{n-1}$.
-On the other hand, one can achieve EMST-ratio of at least $\frac{n-2}{n-1}$ by first computing a minimum spanning tree of the point set in ${O}(n^{2})$ (see [10, 19]), and then coloring the vertices of the two components of it after removing the shortest edge by red and blue.
-Finally, as $\frac{n-2}{n-1}\geq\frac{1}{3}.(3-\frac{4}{n-1})$, this is an ${O}(n^{2})$ time $3$-approximation algorithm.
+On the other hand, one can achieve an EMST-ratio of at least $\frac{n-2}{n-1}$ by first computing a minimum spanning tree of the point set in ${O}(n^{2})$ (see [11, 20]), and then coloring the vertices of the two components of it after removing the shortest edge by red and blue.
+Finally, as $\frac{n-2}{n-1}\geq\frac{(3-\frac{4}{n-1})}{3}$, this is an ${O}(n^{2})$ time $3$-approximation algorithm.
 We remark that a slightly improved running time for computing an EMST is provided in [3].
@@
 Indeed, if we restrict to ${{\mathbb{R}}}^{2}$ (i.e.
-$2$-MAX-EMST-ratio), the approximation factor in Theorem 3.10 can be improved to $2.427$, because the upper bound for the maximum EMST-ratio in the plane is $\approx 2.427$ (see [9]), and for $n>12$ a coloring with EMST-ratio greater than one can be computed in polynomial time (see [13]).
+$2$-MAX-EMST-ratio), the approximation factor in Theorem 3.3 can be improved to $2.427$, because the upper bound for the maximum EMST-ratio in the plane is $\approx 2.427$ (see [10]), and for $n>12$ a coloring with EMST-ratio greater than $1$ can be computed in polynomial time (see [14]).
 As can be observed, the coloring used to present the $3$-approximation is fairly straightforward.
@@
 By a bipartite coloring of a point set, we mean considering one of the Euclidean minimum spanning trees and then partitioning the point set into two subsets labeled by colors in which every edge of the taken EMST is colorful.
-Notice that since the EMST of a point set is not necessarily unique, it may have more than one bipartite colorings, namely one for each EMST.
+Note that since the EMST of a point set is not necessarily unique, it may have more than one bipartite coloring, namely, one for each EMST.
 Intuitively, bipartite colorings give well-distributed colorings of the point set that can be computed in ${O}(n^{2})$ time ($O(n\log n)$ in the plane, see [6]).
-Our experiments suggest that bipartite colorings give a very good approximation of the maximum EMST-ratio for random point sets in the plane (see Section 5).
+Our experiments suggest that bipartite colorings give a very good approximation of the maximum EMST-ratio for random point sets in the plane (see Section 0.A.3).
 Although bipartite colorings seem to be effective, they do not always produce the maximum EMST-ratio.
-Here is an extreme example: Consider $n=2k+1$ points on a zigzag path in the triangular grid and call them a Triangular Chain.
-Figure 4 illustrates a triangular chain for $k=5$.
-The maximum EMST-ratio of a triangular chain with $12k+$ 1 points is $\frac{3k-2}{2k}$ while the ratio for a bipartite coloring is $\frac{2k-1}{2k}$.
-(By stretching the chain slightly, one can ensure a unique EMST and, consequently, a unique bipartite coloring.) As $k$ increases, the proportion between the maximum EMST-ratio and this bipartite EMST-ratio approaches $1.5$.
+Here is an extreme example: Consider $n=2k+1$ points on a zigzag path in the triangular grid and call them a Triangular Chain (see Figure 3).
+By stretching the chain slightly, one can ensure a unique EMST and hence a unique bipartite coloring.
+The maximum EMST-ratio of a triangular chain with $12k+$ 1 points is $\frac{3k-2}{2k}$ (The proof is provided in Section 0.A.1), while the ratio for a bipartite coloring is $\frac{2k-1}{2k}$ As $k$ increases, the proportion between the maximum EMST-ratio and this bipartite EMST-ratio approaches $1.5$.
 Figure 3: An example of a point set where the maximum EMST-ratio has a significant gap from the EMST-ratio of a bipartite coloring.
@@
 We conjecture that this is the worst approximation factor of the maximum EMST-ratio that can be obtained from a bipartite coloring for a point set in the plane.
-Recall that if this conjecture is correct, it would lead to an $O(n\log n)$ time algorithm that achieves a $(1.5)$-approximation for the $2$-MAX-EMST-ratio problem.
-Conjecture 3.12.
+If this conjecture is correct, it would lead to an $O(n\log n)$ time algorithm that achieves a $(1.5)$-approximation for the $2$-MAX-EMST-ratio problem.
+Conjecture 2
 For every point set in ${{\mathbb{R}}}^{2}$, the EMST-ratio for any bipartite coloring is a $(1.5)$-approximation for the maximum EMST-ratio.
-Apart from the bipartite colorings, we are interested in any improved approximation algorithm for the MAX-EXST-ratio that can be obtained by a simple mixed coloring of a point set.
-Question 3.13.
+Apart from the bipartite colorings, we are interested in any improved approximation algorithm for the MAX-EMST-ratio that can be obtained by a simple mixed coloring of a point set.
+Question A
 Does a polynomial-time $2$-approximation algorithm exist for MAX-EMST-ratio?
 4 Average MST-ratio.
-The Average EMST-ratio of a point set $P$ is the average of EMST-ratio over all non-trivial $2$-colorings of $P$.
+The Average EMST-ratio of a point set $P$ is the average of EMST-ratio over all proper $2$-colorings of $P$.
 It is automatically a lower bound for the maximum EMST-ratio.
@@
 for points within a $d$-dimensional unit cube (or any bounded region in ${{\mathbb{R}}}^{d}$).
-Theorem 4.1.
+Theorem 4.1
 For a set of $n$ random points uniformly distributed in $[0,1]^{d}$ the expected value of the average EMST-ratio tends to $d]{2}$ as $n$ goes to infinity.
-Proof 4.2.
+Proof
 Let $n$ be large enough and $P$ be a set of $n$ random points uniformly distributed in $[0,1]^{d}$.
@@
 Ignoring the relatively small $k$ or $n-k$, by (1), we know that for every $k_{0}\leq k\leq n-k_{0}$ we have
-$|EMST(R)|\sim\beta(d).k^{1-1/d},\hskip 7.96677pt|EMST(B)|\sim\beta(d).(n-k)^{1-1/d},\hskip 7.96677pt|EMST(P)|\sim\beta(d).n^{1-1/d}$
-with probability one, where $\beta(d)$ is the constant in (1), and $k_{0}$ is an appropriate constant.
+$|\text{EMST}(R)|\sim$ $\beta(d).k^{1-1/d},\hskip 5.69054pt|\text{EMST}(B)|\sim\beta(d).(n-k)^{1-1/d},$
+$|\text{EMST}(P)|\sim\beta(d).n^{1-1/d}$
+with probability approaching $1$, where $\beta(d)$ is the constant in (1), and $k_{0}$ is an appropriate constant.
 Since $n$ is significantly greater than $k_{0}$, and the EMST-ratio is at most $3$, the contribution of the EMST-ratios for colorings with $k<k_{0}$ red points or with $n-k<k_{0}$ blue points are negligible when computing the average EMST-ratio.
 Thus, the average EMST-ratio over all proper bi-partitions of $P$ is
-$\sum_{P=R\cup B}\frac{|EMST(R)|+|EMST(B)|}{|EMST(P)|}$ $\sim\sum_{k=k_{0}}^{n-k_{0}}\frac{[\beta(d).k^{1-1/d}+\beta(d).(n-k)^{1-1/d}].\binom{n}{k}}{\beta(d).n^{1-1/d}.(2^{n}-2)}$
+$\sum_{P=R\cup B}\frac{|\text{EMST}(R)|+|\text{EMST}(B)|}{|\text{EMST}(P)|}$ $\sim\sum_{k=k_{0}}^{n-k_{0}}\frac{[\beta(d).k^{1-1/d}+\beta(d).(n-k)^{1-1/d}].\binom{n}{k}}{\beta(d).n^{1-1/d}.(2^{n}-2)}$
 $=\sum_{k=k_{0}}^{n-k_{0}}\frac{[k^{1-1/d}+(n-k)^{1-1/d}].\binom{n}{k}}{n^{1-1/d}.(2^{n}-2)}$
-with probability one as $n$ goes to infinity. The right-hand side is a special case of convergence of Bernstein polynomials (see [6, 17]), and therefore converges to $d]{2}$ as $n$ goes to infinity.
-Figure 5: A point set consists of the origin and several points clustered near the three vertices of an equilateral triangle centered at the origin.
-The corresponding minimum spanning tree is depicted.
-As the number of points increases, the average EMST-ratio approaches approximately $2.154$.
+with probability $1$ as $n$ goes to infinity. The right-hand side is a special case of convergence of Bernstein polynomials (see [7, 18]), and therefore converges to $d]{2}$ as $n$ goes to infinity.
+Refer to caption$(a)$Refer to caption$(b)$
+Figure 4: (a) A point set whose average EMST-ratio approaches approximately $2.154$.
+(b) A point set with average EMST-ratio $<1$.
+The point set consists of $n-5$ points placed in a small circle (the core) and $5$ outer points as vertices of a regular pentagon inscribed in the unit circle around the origin.
+The corresponding MST is depicted.
 Certainly, Theorem 3.2 implies that $3$ is an upper bound for the average EMST-ratio of point sets in Euclidean space.
 How much can this upper bound be improved?
-There exist point sets in ${{\mathbb{R}}}^{2}$ with average EMST-ratio $\approx 2.154$.
-For instance, consider the unit circle through the origin in the plane, an equilateral triangle $abc$ on it, and then a point set $P=A\cup B\cup C\cup\{0\}$, where $A$ (resp.
+There exist point sets in ${{\mathbb{R}}}^{2}$ with an average EMST-ratio $\approx 2.154$.
+For instance, consider the unit circle centered at the origin in the plane, an equilateral triangle $abc$ on it, and then a point set $P=A\cup B\cup C\cup\{0\}$, where $A$ (resp.
 $B$, $C$) consists of many points that lie within distance $\varepsilon$ of $a$ (resp.
-$b$,$c$), for $\varepsilon$ being small enough, see Figure 5.
-In this case, the average EMST-ratio approaches $\frac{3+2\sqrt{3}}{3}\approx 2.154$ as the number of points increases, because with high probability, in a random bi-partition, each of the sets $A$, $B$ and $C$ will contain both colors.
-As for the lower bound, we show that the average EMST-ratio for $n$ points is always at least $\frac{n-2}{n-1}$. This is a stronger result than the first part of Theorem 2 in [13].
-Theorem 4.3.
+$b$, $c$), for $\varepsilon$ being small enough (see Figure 4.a).
+In this case, the average EMST-ratio approaches $\frac{3+2\sqrt{3}}{3}\approx 2.154$ as the number of points increases, because with high probability, in a random bi-partition, each of the sets $A$, $B$, and $C$ will contain both colors.
+Regarding the lower bound, we show that the average EMST-ratio for $n$ points is always at least $\frac{n-2}{n-1}$.
+This is a stronger result than the first part of Theorem 2 in [14].
+Theorem 4.2
 For every set of $n$ points in Euclidean space, the average EMST-ratio is at least $\frac{n-2}{n-1}$.
-Proof 4.4.
-Consider a set of $n$ points $P$ and let $e_{1},...,e_{n-1}$ be the edges in a Euclidean MST, $T$, of $P$.
+Proof
+Consider a set of $n$ points $P$ and let $e_{1},...,e_{n-1}$ be the edges of an EMST of $P$, called $T$.
 Assume that the edge $e_{i}$ has weight $w_{i}$ and $w_{1}\leq w_{2}\leq\ldots\leq w_{n-1}$.
 Considering a coloring $P=R\cup B$ of $P$ into red and blue points, let $i$ be the smallest index such that the edge $e_{i}$ is colorful.
-$EMST(R)\cup EMST(B)\cup\{e_{i}\}$ is a spanning tree of $P$ and therefore, its weight is at least $|EMST(P)|$.
-This implies $|EMST(R)|+|EMST(B)|\geq|EMST(P)|-w_{i}$.
-On the other hand, considering a random coloring of $P$ with two colors, the probability for the edge $e_{i}$ being colorful is $\frac{1}{2}$, and since the $e_{i}$’s do not form a cycle, their corresponding events are independent.
+As $\text{EMST}(R)\cup\text{EMST}(B)\cup\{e_{i}\}$ is a spanning tree of $P$, it holds that $|\text{EMST}(R)|+|\text{EMST}(B)|\geq|\text{EMST}(P)|-w_{i}$.
+On the other hand, considering a random coloring of $P$ with two colors, the probability that the edge $e_{i}$ is colorful is $\frac{1}{2}$, and since the $e_{i}$’s do not form a cycle, their corresponding events are independent.
 Thus, for a fixed index $j$, the probability that in a random coloring, the shortest colorful edge of $T$ is $e_{j}$ is $(\frac{1}{2})^{j}$.
 Therefore, there are precisely $2^{n-j}$ colorings of $P$ in which $j$ is the smallest index such that the edge $e_{j}$ is colorful.
-By the above argument and since there are $22^{n}-$ 2 possible (proper) colorings, the average EMST-ratio is at least:
-$|EMST(P)|-\sum_{i=1}^{n-1}\frac{2^{n-i}w_{i}}{2^{n}-2}.$
-As $\sum_{i=1}^{n-1}2^{n-i}=2^{n}-2$ and $w_{1}\leq w_{2}\leq\ldots\leq w_{n-1}$, then:
-$|EMST(P)|-\sum_{i=1}^{n-1}\frac{2^{n-i}w_{i}}{2^{n}-2}\geq|EMST(P)|-\frac{|EMST(P)|}{n-1},$
-and hence the claim.
+By the above argument and since there are $22^{n}-$ 2 possible (proper) colorings, the average EMST-ratio $a$ is at least:
+$a\geq|\text{EMST}(P)|-\sum_{i=1}^{n-1}\frac{2^{n-i}w_{i}}{2^{n}-2}.$
+As we have $\sum_{i=1}^{n-1}2^{n-i}=2^{n}-2$ and $w_{1}\leq w_{2}\leq\ldots\leq w_{n-1}$, it can be concluded that $a\geq|\text{EMST}(P)|-\frac{1}{n-1}\cdot|\text{EMST}(P)|$, and hence the claim.
 Note that Theorem 4.2 holds also for general graphs.
@@
 The only instances we found in ${{\mathbb{R}}}^{2}$ consist of point sets where $n-6\leq k\leq n-1$ points are clustered closely together, forming a core, and $1\leq n-k\leq 6$ points are sufficiently far from the core such that their pairwise distances are not less than their distances to the core.
-We believe those are the only point sets with an average EMST-ratio of less than $1$.
-As an example, we show that the average EMST-ratio for the point set in Figure 6 is less than $1$.
+We believe these are the only point sets with an average EMST-ratio of less than $1$.
+As an example, in Section 0.A.2, we show that the average EMST-ratio for the point set in Figure 4.b is less than $1$.
+Question B
+Which point sets in ${{\mathbb{R}}}^{2}$ have an average EMST-ratio $<1$?
+5 Discussion
+In addition to the conjectures and questions posed in this paper, there are several other directions that are worth exploring further.
+Coloring with Many Crossings.
+For a point set $P$ in the plane, partitioned as $P=R\cup B$ of $P$, the Chromatic Crossing Number is the number of crossings between pairs of edges, one in $\text{EMST}(R)$ and one in $\text{EMST}(B)$.
+Intuitively, this also measures how mixed the coloring is; the higher the number, the more mingled it appears.
+In fact, for any point set, it is easy to find colorings with no crossings between the two MSTs.
+More interesting, however, is the Maximum Chromatic Crossing Number, which represents the highest chromatic crossing number achievable across all possible colorings of the point set.
+Figure 5: A point set and a bi-partition with chromatic crossing number $4$.
+Question C
+How large and how small can the maximum chromatic crossing number of a set of $n$ points in the plane be?
+Question D
+Are there relations between the chromatic crossing number and the EMST-ratio? In particular, does the coloring that maximizes the chromatic crossing number tend to have a relatively high EMST-ratio?
+Conjecture 3
+Finding the maximum chromatic crossing number of a point set in the plane is NP-hard.
+Maximum MST Length over Subsets.
+A related question to finding the maximum MST-ratio is the MAX-EMST-subset problem.
+Here, given a point set $P$, we aim to find the maximum EMST length over all subsets $Q\subseteq P$.
+This extends to the abstract setting as the MAX-MST-subset problem, where, given a weighted graph $G$, we seek the maximum MST weight over all induced subgraphs $H\subseteq G$.
+Question E
+Are MAX-MST-subset and MAX-EMST-subset NP-hard?
+More Colors.
+Similar questions can be asked for point sets with $c>2$ colors.
+To express the MST-ratio, we consider the total length of MST of points in each color divided by the length of an MST of the (uncolored) point set.
+Theorem 3.2 can be generalized to obtain $12c-$ 1 as an upper bound for the MST-ratio in any metric space.
+For points in Euclidean space, this bound can be improved for $c\geq 4$.
+In particular, as $\rho_{d}\geq 0.577$ for every $d$, the maximum EMST-ratio is upper bounded by $\frac{c}{0.577}$.
+It remains unclear how the results in [10] on lattices can be generalized to more than $2$ colors.
+Furthermore, exploring the average EMST-ratio for $c$ colors is an interesting direction to pursue.
+Appendix 0.A Appendix
+0.A.1 Maximum EMST-ratio of Triangular Chain
+In this section, we show that the maximum EMST-ratio for the triangular chain, shown in Figure 3 is $\frac{3k-2}{2k+1}$.
+Let $P$ be the point set of a triangular chain with $12k+$ 1 points and let us number the points of $P$ from left to right by $p_{1}$, $p_{2}$, $\ldots$, $p_{2k+1}$.
+As $|\text{EMST}(P)|=2k$, to prove the claim we show that for any proper coloring of the chain to red and blue, we have $|\text{EMST}(R)|+|\text{EMST}(B)|\leq 3k-2$.
+We prove this by applying induction on $k$.
+The claim holds trivially for $k=1$, as any proper coloring has $|\text{EMST}(R)|+|\text{EMST}(B)|=1=3-2$.
+Now, assume that the claim holds for $k$.
+Now we prove the claim for $k+1$.
+Assume we have a chain with $12(k+1)+$ 1 points.
+Let $B$ and $R$ be the set of blue points and red points in a proper coloring of $P$.
+We distinguish two cases:
+a) If there exists $i\in\{1,3\}$, such that $p_{i}$, $p_{i+1}$ and $p_{i+2}$ have the same color: W.l.o.g.
+we assume these points belong to $R$.
+First, consider the case $i=1$.
+In this case, $R\setminus\{p_{1},p_{2}\}$ and $B$ is a proper coloring for $P\setminus\{p_{1},p_{2}\}$.
+Also, $T_{R}:=\text{EMST}(R\setminus\{p_{1},p_{2}\})\cup\{p_{1}p_{2},p_{2}p_{3}\}$ is a spanning tree of $R$.
+By induction hypothesis $|\text{EMST}(R\setminus\{p_{1},p_{2}\})|+|\text{EMST}(B)|\leq 3k-2$, which implies that:
+$|\text{EMST}(R)|+|\text{EMST}(B)|\leq|T_{R}|+|\text{EMST}(B)|$
+$\leq 3k-2+|p_{1}p_{2}|+|p_{2}p_{3}|=3k-1+2<3(k+1)-2,$
+and hence the claim.
+Now consider the case $i\neq 1$ (i.e.
+$i=3$).
+In this case, let $P^{\prime}$ be the triangular chain obtained from $P$ by removing $p_{1}$ and $p_{2}$.
+Let $R^{\prime}$ and $B^{\prime}$ be a proper coloring of $P^{\prime}$, where:
+  • For every $j\geq 5$, $p_{j}\in R^{\prime}$ iff $p_{j}\in R$.
+  • For every $j\in\{3,4\}$, $p_{j}\in R^{\prime}$ iff $p_{j-2}\in R$.
+Note that $R^{\prime}$ and $B^{\prime}$ is a proper coloring of $P^{\prime}$ and hence by induction hypothesis:
+$|\text{EMST}(R^{\prime})|+|\text{EMST}(B^{\prime})|\leq 3k-2.$
+As $i\neq 1$, then $(p_{1}\cup p_{2})\cap B\neq\emptyset.$ We have $|\text{EMST}(R^{\prime})|=|\text{EMST}(R)|-2$ and $|\text{EMST}(B^{\prime})|+1\geq|\text{EMST}(B)|$. Therefore:
+$|\text{EMST}(R)+|\text{EMST}(B)|\leq|\text{EMST}(R^{\prime})|+|\text{EMST}(B^{\prime})|+3\leq 3(k+1)-2.$
+b) Otherwise: W.l.o.g.
+assume $p_{3}\in R$.
+First, assume $p_{1}$ and $p_{2}$ are in $B$.
+Since we are not in case a), then at least one of the points $p_{4}$ and $p_{5}$ is in $B$.
+Therefore, $R$ and $B\setminus\{p_{1},p_{2}\}$ is a proper coloring for $P\setminus\{p_{1},p_{2}\}$.
+Now $T_{B}:=\text{EMST}(B\setminus\{p_{1},p_{2}\})\cup\{p_{1}p_{2},p_{2}p_{i}\}$ is a spanning tree of $B$, where $p_{i}\in\{p_{4},p_{5}\}\cap B$.
+Note that by the induction hypothesis $|\text{EMST}(B\setminus\{p_{1},p_{2}\})|+|\text{EMST}(R)|\leq 3k-2$, therefore
+$|\text{EMST}(R)|+|\text{EMST}(B)|\leq|T_{B}|+|\text{EMST}(R)|$
+$\leq 3k-2+|p_{1}p_{2}|+|p_{2}p_{i}|=3k-1+|p_{2}p_{i}|<3(k+1)-2.$
+Now assume that there exists $j\in\{1,2\}$ such that $p_{j}\in R$.
+Since we are not in case a), then $p_{3-j}$ is blue and at least one of the points $p_{4}$ and $p_{5}$ (say $p_{i}$) is in $B$.
+Therefore, $R$ and $B\setminus\{p_{1},p_{2}\}$ is a proper coloring for $P\setminus\{p_{1},p_{2}\}$.
+Now $T_{B}:=\text{EMST}(B\setminus\{p_{1},p_{2}\})\cup\{p_{j}p_{3},p_{3-j}p_{i}\}$ is a spanning tree of $B$.
+Note that by induction hypothesis $|\text{EMST}(B\setminus\{p_{1},p_{2}\})|+|\text{EMST}(R)|\leq 3k-2$, therefore
+$|\text{EMST}(R)|+|\text{EMST}(B)|\leq|T_{B}|+|\text{EMST}(R)|$
+$\leq 3k-2+|p_{j}p_{3}|+|p_{3-j}p_{i}|=3k-1+|p_{3-j}p_{i}|\leq 3(k+1)-2.$
+0.A.2 A point set with average EMST-ratio $<1$
+In this section, we show that the average EMST-ratio for the point set described in Figure 4.b is smaller than $1$.
 The point set consists of $n-5$ points in a small circle through the origin whose radius is $\varepsilon\in{{o}}(\frac{1}{n2^{n}})$ (the core), along with $5$ outer points as vertices of a regular pentagon inscribed in the unit circle around the origin.
-In this example, clearly $|EMST(P)|\geq 5-{{o}}(n\varepsilon)$.
+Observe that each side of this pentagon is $s=2\sin{36}\approx 1.176$ and each diagonal is $d=\frac{1+\sqrt{5}}{2}s\approx 1.902$.
+In this example, clearly $|\text{EMST}(P)|\geq 5-{O}(n\varepsilon)$.
 In any bi-coloring of $P$ by red and blue if an outer point has the same color as one of the core points, then the edge connecting this outer point to a core point appears in one of the monochromatic EMSTs.
-Therefore, in most of the bi-colorings of $P$ where the core contains both colors, we have $|EMST(R)|+|EMST(B)|\leq 5+{{o}}(n\varepsilon)$.
-There are only $62$ many proper bi-colorings with a mono-chromatic core, and one can show that in each of them $|EMST(R)|+|EMST(B)|<5.1$.
-In fact, in $52$ of such colorings we even have $|EMST(R)|+|EMST(B)|<4.8$.
-Thus the average EMST-ratio for this point set is upper bounded by:
-$\frac{(2^{n}-64)\times(5+{{o}}(n\varepsilon))+52\times(4.8)+10\times(5.1)}{(2^{n}-2)\times(5-{{o}}(n\varepsilon))}<1.$
-Figure 6: A point set with average EMST-ratio smaller than $1$.
-The point set consists of $n-5$ points in a very small circle through the origin (the core), and $5$ outer points as vertices of a regular pentagon inscribed in the unit circle around the origin.
-The corresponding MST is depicted.
-Question 4.5.
-Can the point sets in ${{\mathbb{R}}}^{2}$ with an average EMST-ratio $<1$ be characterized?
-5 Computational Experiments
+Therefore, in $642^{n}-$ 64 bi-colorings of $P$ where the core contains both colors, we have $|\text{EMST}(R)|+|\text{EMST}(B)|\leq 5+{O}(n\varepsilon)$.
+There are only $622^{6}-2=$ 62 proper bi-colorings with a mono-chromatic core.
+Consider one of these $62$ colorings and assume the points in the core are red.
+In this case, if $50\leq n_{r}<$ 5 denotes the number of outer points that are red, then $|\text{EMST}(R)|=n_{r}+O(n\varepsilon)$.
+Also, if the outer blue points are consecutive vertices of the outer pentagon, $|\text{EMST}(B)|=(4-n_{r})s$ and otherwise $|\text{EMST}(B)|=(3-n_{r})s+d$.
+So in each of these colorings $|\text{EMST}(R)|+|\text{EMST}(B)|\leq d+s+2+{O}(n\varepsilon)<5.1$.
+Moreover, in $62-2\times 5=52$ of such colorings we even have $|\text{EMST}(R)|+|\text{EMST}(B)|\leq d+3+{O}(n\varepsilon)<4.91$.
+Thus, the average EMST-ratio for this point set is upper bounded by:
+$\frac{(2^{n}-64)\cdot(5+{O}(n\varepsilon))+52\cdot(4.91)+10\cdot(5.1)}{(2^{n}-2)\cdot(5-{O}(n\varepsilon))}<1.$
+0.A.3 Computational Experiments
 This section presents the results of computational experiments comparing the average EMST-ratio, maximum EMST-ratio, and the EMST-ratio for a bipartite coloring, all conducted on random point sets.
@@
 Figure 6: The mean values of the maximum EMST-ratios (green), the bipartite EMST-ratios (blue), and the average EMST-ratios (red) of $500$ random sets of $n$ points, where $5\leq n\leq 20$.
-Question 5.1.
+Question F
 What is the limit for the expected maximum EMST-ratio of $n$ points sampled uniformly at random in $[0,1]^{2}$, as $n$ goes to infinity?
-Conjecture 5.2.
+Conjecture 4
 For $n$ random points uniformly distributed in $[0,1]^{2}$, the maximum EMST-ratio is less than $2$ with probability tending to $1$ as $n\rightarrow{\infty}$.
@@
 Figure 7: Each point in the diagram represents a random point cloud, with the $x$ and $y$ coordinates corresponding to a bipartite EMST-ratio of the sample point set computed using Prim’s algorithm and the maximum EMST-ratio, respectively.
-Conjecture 5.3.
+Conjecture 5
 For $n$ random points uniformly distributed in $[0,1]^{2}$, the maximum EMST-ratio is less than $1.1$ times any bipartite EMST-ratio of the point set with probability tending to $1$ as $n$ goes to infinity.
@@
 This suggests that bipartite colorings of a point set provide better EMST-ratios than what is expected from a random coloring.
-Although this is not always the case for finite point sets, we believe there are very few counterexamples.
-Question 5.4.
+Although this is not always the case for finite point sets, we believe that there are very few counterexamples.
+Question G
 Can the point sets in ${{\mathbb{R}}}^{2}$ for which a bipartite EMST-ratio of them is smaller than the average EMST-ratio be characterized?
-6 Discussion
-In addition to the conjectures and questions posed in this paper, there are several other directions worth exploring further.
-Coloring with Many Crossings.
-For a point set $P$ in the plane and a partition $P=R\cup B$ of $P$ into red and blue points, the Chromatic Crossing Number is defined as the number of crossings between pairs of edges, one in $EMST(R)$ and one in $EMST(B)$.
-Intuitively, this number can be viewed as another measure of how mixed the coloring is; the higher the chromatic crossing number, the more mingled the coloring appears.
-In fact, for any point set, it is easy to find colorings with no crossings between the two MSTs.
-More interesting, however, is the Maximum Chromatic Crossing Number, which represents the highest chromatic crossing number achievable across all possible colorings of the point set.
-Figure 10: An example of a point set and a bi-partition with chromatic crossing number $4$.
-Question 6.1.
-How large and how small can the maximum chromatic crossing number of a set of $n$ points in the plane be?
-Question 6.2.
-Are there relations between the chromatic crossing number and the EMST-ratio? In particular, does the coloring that maximizes the chromatic crossing number tend to have a relatively high EMST-ratio?
-Conjecture 6.3.
-Finding the maximum chromatic crossing number of a point set in the plane is NP-hard.
-Maximum MST Length over Subsets.
-Related question to finding the maximum MST-ratio is the following.
-In the MAX-EMST-subset problem, given a point set $P$, we aim to find the maximum EMST length over all subsets $Q\subseteq P$.
-The question can be raised also in the abstract setting.
-Namely, in the MAX-MST-subset problem, given a weighted graph $G$, we would like to find the maximum MST weight over all induced subgraphs $H\subseteq G$.
-Question 6.4.
-Are the problems MAX-MST-subset and MAX-EMST-subset NP-hard?
-More Colors.
-Similar questions can be asked for point sets with $c>2$ colors.
-To express the MST-ratio, we consider the total length of minimum spanning trees of points in each color divided by the length of a minimum spanning tree of the (uncolored) point set.
-Theorem 3.8 can be generalized to obtain $12c-$ 1 as an upper bound for the MST-ratio in any metric space.
-For points in Euclidean space, this bound can be improved for $c\geq 4$.
-In particular, as $\rho_{d}\geq 0.577$ for every $d$, the maximum EMST-ratio is upper bounded by $\frac{c}{0.577}$.
-It remains unclear how the results in [9] on lattices can be generalized to more than $2$ colors.
-Additionally, exploring the average EMST-ratio for $c$ colors is an interesting direction to pursue.
-Acknowledgements. The authors thank Ondřej Draganov for his valuable insights on the topic of average EMST-ratio.